### v1

# Analytic Continual Test-Time Adaptation for Multi-Modality Corruption

Abstract
Test-Time Adaptation (TTA) aims to help pre-trained model bridge the gap between source and target datasets using only the pre-trained model and unlabelled test data. A key objective of TTA is to address domain shifts in test data caused by corruption, such as weather changes, noise, or sensor malfunctions. Multi-Modal Continual Test-Time Adaptation (MM-CTTA), an extension of TTA with better real-world applications, further allows pre-trained models to handle multi-modal inputs and adapt to continuously-changing target domains. MM-CTTA typically faces challenges including error accumulation, catastrophic forgetting, and reliability bias, with few existing approaches effectively addressing these issues in multi-modal corruption scenarios. In this paper, we propose a novel approach, Multi-modality Dynamic Analytic Adapter (MDAA), for MM-CTTA tasks. We innovatively introduce analytic learning into TTA, using the Analytic Classifiers (ACs) to prevent model forgetting. Additionally, we develop Dynamic Selection Mechanism (DSM) and Soft Pseudo-label Strategy (SPS), which enable MDAA to dynamically filter reliable samples and integrate information from different modalities. Extensive experiments demonstrate that MDAA achieves state-of-the-art performance on MM-CTTA tasks while ensuring reliable model adaptation.
1 Introduction
Test-Time Adaptation (TTA) aims to help the pre-trained model bridge the gap between the source domain and the target domain (Wang et al., 2021; Liang et al., 2024). Unlike Unsupervised Domain Adaptation (UDA) (Zhang et al., 2015; Liang et al., 2024), TTA performs adaptation without the need for any source data (i.e., pre-trained dataset), which not only saves computational resources by avoiding retraining but also preserves the privacy of the source data. One key TTA application is addressing the problem of domain shift from source data to corrupted test data, where the corruption is often caused by external factors (e.g., weather changes, ambient noise) or sensor malfunctions. As an extension of TTA, Continual Test-Time Adaptation (CTTA) has been proposed to align with real-world scenarios where domain shifts usually are dynamic (Wang et al., 2022). Challenges in CTTA mainly consist of error accumulation and catastrophic forgetting. Error accumulation, stemming from incorrect pseudo-labels, can mislead models’ adaptation and potentially lead to collapse (Chen et al., 2019). Catastrophic forgetting refers to the loss of knowledge from the source data during continuous adaptation, reducing the model’s generalization ability (McCloskey & Cohen, 1989). To address these challenges, various CTTA methods have been proposed, yielding promising results in corruption-related tasks (Wang et al., 2022; Gao et al., 2023; Niu et al., 2022).
Most existing CTTA approaches focus solely on single-modal scenarios, paying less attention to multi-modal applications. Compared with TTA and CTTA, Multi-Modal Continual Test-Time Adaptation (MM-CTTA) (Cao et al., 2023) shows greater potential for real-world applications, as multi-modal data integrates a broader range of information than single-modality adaptation, resulting in more robust networks (Radford et al., 2021). However, applying existing CTTA methods to MM-TTA by simply replacing the backbone with muiti-modal encoders is less optimal. MM-CTTA performance can easily suffer from reliability bias, where intra-modality domain shifts increase information discrepancies in downstream fusion networks (Yang et al., 2024). Such effect becomes more pronounced when modality corruption changes dynamically. Although a few works address Multi-Modal Test-Time Adaptation (MM-TTA) (Yang et al., 2024; Lei & Pernkopf, 2024), they struggle with model forgetting during continuous adaptation and the challenges posed by interleaved modality corruption.
To further illustrate, we use a representative TTA method, TENT (Wang et al., 2021), as an example in Fig.1. The results indicate that traditional TTA methods face significant limitations due to error accumulation, catastrophic forgetting, and reliability bias. These challenges suggest that methods designed for MM-CTTA must meet three key requirements: (1) effectively mitigating error accumulation during adaptation, (2) retaining knowledge of the source data after adapting to various domains, and (3) dynamically suppressing the influence of corrupted modalities while prioritizing more reliable ones. To the best of our knowledge, the method closest to meeting these criteria is CoMAC (Cao et al., 2023), which focuses on segmentation tasks in living environments rather than addressing modality corruption.
In this paper, we propose a new approach named the Multi-modality Dynamic Analytic Adaptor (MDAA) to address the challenges in MM-CTTA. MDAA comprises three primary components: (i) the Analytic Classifiers (ACs), (ii) the Dynamic Selection Mechanism (DSM), and (iii) the Soft Pseudo-label Strategy (SPS). ACs update the model by addressing a recursive ridge regression problem, optimizing on both new target data and learned knowledge to avoid catastrophic forgetting. DSM selectively updates each AC based on its output reliability, thereby alleviating reliability bias. The SPS enhances the model’s robustness to label noise by assigning varying probabilities to multiple labels, which mitigates potential error accumulation. contribution of thiss work can be summarised as follows:
1. We c
1). We propose a method, MDAA, for a more challenging TTA task named MM-CTTA, and explain why typical TTA methods are not well-suited for this task, as illustrated by a example in Fig.1.
2). We innovatively apply AC to TTA to keep the model from catastrophic forgetting during adaption. We propose DSM and SPS to further dynamically integrate the features of different modal thus mitigating reliability bias the error accumulation.
3). We design two MM-CTTA tasks to meet the needs of real-world environments. Extensive experiments demonstrate that MDAA achieves SOTA performance, surpassing previous methods by up to 6.22% and 6.84% in the two tasks respectively.
2 Related Works
2.1 Test-Time Adaptation
Test-Time Adaptation (TTA) focuses on enabling a pre-trained model to adapt to a new target domain without requiring access to the source domain data used to initially train the model. A major challenge in TTA is error accumulation, which affects methods that rely on pseudo-labeling; incorrect predictions can mislead the adaptation process, potentially causing the model to collapse (Chen et al., 2019). One of the early solutions, TENT, addresses this by only updating the model’s batch normalization (BN) layers through entropy minimization (Wang et al., 2021). Subsequent research (Niu et al., 2023; Gong et al., 2022a; Zhou et al., 2023) has further mitigated this issue by filtering out low-confidence predictions using carefully designed thresholds and updating the layer normalization (LN) layers for more robust performance.
2.2 Continual Test-Time Adaptation
Continual Test-Time Adaptation (CTTA) extends TTA to scenarios where the target domain changes continuously in an online manner without access to source data. The additional challenge CTTA faces is known as catastrophic forgetting, which occurs when a model adapts to the target domain, leading to the loss of knowledge acquired from the source domain and dimension of generalization ability. To solve this problem, some studies turn to Continual Learning (CL) and achieve great success (Niu et al., 2022; Cao et al., 2023). We follow this trend and introduce a novel CL approach called Analytic Continual Learning (ACL) in MDAA.
ACL provides a global optimal solution through matrix inverse operations (Guo & Lyu, 2004). To address the out-of-memory issue caused by large inverse matrices, Zhuang et al. (2021a) demonstrated that iterative computation using block-wise data achieves results equivalent to joint computation, making analytic learning highly effective in continual learning. By treating data from different time periods as blocks, ACL allows for recursive computation, with the final result being as accurate as if all data were processed simultaneously. Thanks to its non-forgetting properties, ACL has shown strong performance across various CL tasks in recent years (Zhuang et al., 2022; 2023). Inspired by ACL’s success in CL tasks, we apply ACL to TTA for the first time in this work, implementing several enhancements to address the issue of catastrophic forgetting.
2.3 Multi-modality Test-Time Adaptation
Multi-modality Test-Time Adaptation (MM-TTA) seeks to enhance model reliability by incorporating multi-modal data into the TTA task. However, imbalances in inter-modal reliability can result in significant performance degradation, known as reliability bias (Wang et al., 2020). Most existing MM-TTA models address this issue by independently updating BN or LN layers of each feature encoder, followed by a weighted fusion mechanisms (Shin et al., 2022; Cao et al., 2023; Xiong et al., 2024). Although this allows more reliable modalities to carry more weight during fusion, the approach remains relatively shallow in terms of information integration.
A recent model called READ (Yang et al., 2024) introduced a more advanced approach by fusing features through a Vision Transformer (ViT) (Dosovitskiy et al., 2020) block, which allows for the preservation of parameters inherited from source data while effectively integrating inter-modal information (Vaswani, 2017; Gong et al., 2022b). READ achieves reliable adaptation by modulating only the fusion layer within the attention module of the ViT block. Follow-up work (Lei & Pernkopf, 2024) aimed to improve performance by updating both the feature encoders and the fusion layer. However, this approach requires prior knowledge about which modalities are corrupted, making it less capable for real-world scenario.
In this paper, our MDAA approach introduces classifiers for each feature encoder and the fusion network. By adjusting the parameters of each classifier individually, MDAA maximizes the use of information from each modality, thus mitigating the effects of reliability bias. Crucially, these classifier updates for upstream and downstream blocks do not cause conflicts, as the entire pre-trained model remains frozen throughout the process. This property enables MDAA to adapt to changing modality corruption without extra prior knowledge, distinguishing it from existing MM-TTA approaches.
3 Method
In this section, we first define the challenging MM-CTTA setting and introduce notations for key concepts in Sec. 3.1. We then introduce the proposed Multi-modality Dynamic Analytic Adapter (MDAA) approach, which includes Analytic Classifiers (ACs) integrated with pre-trained multi-modal encoders (Sec. 3.2), the Dynamic Selection Mechanism (DSM) (Sec. 3.3), and the Soft Pseudo-label Strategy (SPS) (Sec. 3.4). An overview of MDAA is illustrated in Fig. 2, and the pseudo-code for MDAA is provided in Appendix B.
3.1 Problem Definition and Notations
In this paper, we focus on an audio-video classification task as an example, using two modalities for illustration without loss of generality. In MM-CTTA, the pre-trained model is trained on a labeled source dataset in source domain S, where and represent the audio and video training data matrices respectively. represents the corresponding one-hot label set. During the adaptation phase, for each timestamp in the target domain T, the model must perform inference and update its parameters based on the unlabeled test dataset . The suffix indicates the current target domain, as it shifts continually. Note that only the test dataset at timestamp is available for updating the parameters .
3.2 Source Model and Analytic Classifier Configuration
In the context of multi-modality classification task, we propose to integrate the ACs as classifiers into a typical extraction-fusion approach. The standard structure can be represented as follows:
where and represent upstream feature encoders for audio and video modalities, respectively, indicates the fusion operation implemented as concatenation in this paper, denotes the downstream fusion network and indicates the classifier.
Specifically, we leverage multiple ACs as classifiers for extracted features of each modality and the fused features, with each classifier making an independent prediction. Each AC is of the same structure as a two-layer fully connected network, denoted by . Features are first non-linearly projected into a higher dimensional space by the feature expansion layer , enhancing their expressiveness (Zhuang et al., 2022; 2021b). The projected features are then passed through the linear layer for classification. The feature expansion layer remains frozen, while the linear layer requires updates during the adaptation process. For simplicity, we combine the feature encoder and the feature expansion layer into a single function, denoted as , and refer to the projected features as . Consequently, MDAA yields three classifiers as follows:
Since all ACs can be represented by the same equation, we will not distinguish between them except for special needs in the following discussion. Following previous ACL works Zhuang et al. (2022), the classifier are updated by solving the ridge regression to optimize on the source data . To further solve the class imbalance problem within the source data, inspired by Fang et al. (2024), we formulate the optimization problem as follows:
where indicates the Frobenius norm, is the regularization parameter and is the samples size of . , and represents the weight, expanded feature vector and one-hot label of sample k in , while the weight is further defined as
where is the number of classes in and is the number of samples in from category to which sample belongs. Following the ridge regression solution, the solution to optimization problem is given in Theorem 1.
Theorem 1. The optimal solution to Formula 4 is given as
where and . The proof of Theorem 1 is provided in Appendix A. In addition to the classifier weights , a memory bank needs to be constructed during the training phase. Unlike other methods (Cao et al., 2023; Zhang et al., 2023; Xiong et al., 2024), our memory bank contains only two types of matrices, which can be represented as , where
Both and are used to extract and preserve the learned knowledge from the source dataset, which cannot be accessed during adaptation. Therefore weight can be further rewritten as
3.3 Dynamic Selection Mechanism
When adapting to sequentially incoming multi-modal data in MM-CTTA, some modality may be unreliable due to corruption. Such corrupted data can mislead the model to learn incorrect information. To address the reliability bias, we propose Dynamic Selection Mechanism (DSM) to determine whether each AC should be updated in a dynamic way. M first identifies the most reliable classifier among the three (i.e., a, v, m) as the leader, while the other classifiers are treated as follower.
Specifically, the leader is determined by comparing the maximum probability from each classifier’s distribution. For each sample, the model prediction corresponds to the leader’s prediction, and the leader will not be updated to maintain class balance. Whether to update each follower in the model depends on the comparison between its maximum probability distribution and the leader’s maximum probability distribution . We consider four possible scenarios, as illustrated in Fig.3:
(i). Close Distributions: and are quite close and refer to the same label. In this case, the follower is not updated, as this would only reinforce what it has already learned, potentially leading to an imbalanced class distribution.
(ii). Different Labels with Close Probabilities: and are quite close while they refer to different labels. The follower is also not updated in this scenario because there is no certainty that the leader’s result is correct, and updating may introduce errors.
(iii). Evenly Distributed Probabilities: Both the leader and follower have evenly distributed probabilities with no significant difference between labels. Again, no update occurs.
(iv). Significant Difference: The leader has a higher probability label, while the follower has a more even probability distribution. In this scenario, the follower should be updated.
In general, cases (i),(ii) and (iii) all belong to a small gap between and , while case (iv) belongs to a larger gap. Therefore, given a pre-defined threshold , DSM can be noted as:
3.4 Soft Pseudo-label Strategy
Using soft labels in self-supervised learning is quite popular as it preserves a broader range of possible outcomes compared to hard (one-hot) label learning (Müller et al., 2019; Hinton et al., 2015). Inspired by such trend, we use soft pseudo-labels to update the ACs during MM-CTTA, as illustrated in Fig.2(C). For each test sample , we choose the top classes of leader’s distribution set and assign them with weights , () respectively. The reconstructed label for sample through SPS can be represented as
Since ACL considers global optimization, which means it accounts not only for the input data at timestamp , but also all previous data processed by the model, including and . With the reconstructed label throughout , the optimization problem for weight matrix can be represented as:
It is important to note that while we treat source data and target data separately, we adjust the category balance for the source dataset by assigning weights to each class. In contrast, it is challenging to apply this approach to the target dataset, as it is impossible to know the exact number of samples in each category within the target domain. However, with the help of the DSM, category balance can still be maintained, as unimportant samples have been filtered out. Given that the weights added to the sample data are expected to average out to 1, the impact of each sample from both the source and target domains on the model can be considered equivalent.
Due to the definition of TTA, all datasets (i.e., and ) prior to timestamp are not accessible when solving the optimization problem. However, with the aid of the memory bank , the solution can still be computed, as stated in Theorem 2.
Theorem 2. The optimal solution to Formula 12 is given as
where the memory bank is updated in a recursive way with timestamp , as
The proof of Theorem 2 is also provided in Appendix A. It can be seen that, the size of the memory bank is constant and depends on the dimension of the feature expansion layer and class number . The equations above demonstrate that the size of remains unchanged during adaptation, regardless of the timestamp .
4 Experiments
In this section, we evaluate the proposed MDAA approach under the challenging MM-CTTA setting. The MM-CTTA setting is detailed in Sec. 4.1. Sec. 4.2 compares MDAA with SOTA methods through extensive experiments to demonstrate its superior performance. Additionally, the ablation studies on each component of MDAA are presented in Sec. 4.3, which illustrate the effectiveness of MDAA in addressing different challenges within the MM-CTTA setting. The implementation details are provided in Appendix C.
4.1 Benchmarks and Settings
In this section, we introduce the datasets and task settings used for MM-CTTA. The MM-CTTA setting requires the model to initially train on uncorrupted source datasets. Subsequently, the model performs TTA on each corrupted target domain in sequence. We utilize two datasets for this setting: Kinetics50 (Kay et al., 2017) and VGGSound (Chen et al., 2020). While the original uncorrupted datasets serve as source, the corrupted target datasets, Kinetics50-C and VGGSound-C, are constructed following Yang et al. (2024), which introduces 15 types of video corruptions and 6 audio corruptions at severity level 5.
To evaluate the model performance, we designed two classification tasks specifically for MM-CTTA following previous research on corruption-related TTA (Wang et al., 2022; 2021; Yang et al., 2024; Lei & Pernkopf, 2024). The first task, named progressive single-modality corruption, sequentially introduces different types of corruption to one modality while keeping the other modality uncorrupted. Focusing on evaluating the model’s resistance to catastrophic forgetting, this task is set in an online manner, where the model processes only one sample at a time. The second task, called interleaved modality corruption, continually alternates corruption between the two modalities. While most methods perform poorly in the online setting due to severe catastrophic forgetting, this task uses a batch size of 64 during test time to emphasize assessing the model’s ability to adapt to dynamic reliability biases.
4.2 Performance Comparison
To provide a comprehensive comparison, we reproduce different types of TTA methods under the MM-CTTA setting. Typical TTA methods of TENT (Wang et al., 2021) and SAR (Niu et al., 2023); CTTA methods of CoTTA (Wang et al., 2022) and EATA (Niu et al., 2022); and MM-TTA methods of MMTTA (Shin et al., 2022) and READ (Xiong et al., 2024). To ensure a fair comparison, all methods are based on the pre-trained ViT-baesed CAV-MAE (Gong et al., 2022b) as the multi-modal encoders. When reproduce methods that update the BN layers, we instead update the LN layers to suit the ViT structure. Additionally, we evaluate the Source model as a strong baseline, which is only trained on the source dataset and remains frozen during test time.
Performances of different methods on progressive single-modality corruption are listed in Table 1, 2 and 3, where Table 1 is based on the audio corruption, Table 2 and 3are performance on the video corruption in Kinetics50-C and VGGSound-C. The proposed MDAA achieves SOTA performance on most target domains. It is noteworthy that nearly all comparison models collapse and perform significantly worse than non-updating models on VGGSound-C. In contrast, EATA achieves better results by restricting model parameter updates, which facilitates successful adaptation. Specifically, MDAA outperforms previous methods by 3.00%-3.57% and 3.03%-6.22% on average for audio and video tasks in Kinetics50-C, and by 0.94%-1.10% and 0.13%-0.18% on average for audio and video tasks in VGGSound-C. Furthermore, MDAA consistently maintains its lead over other methods in the later stages of the tasks. These results demonstrate that MDAA is more robust against catastrophic forgetting in MM-CTTA.
| Method | Type | Kinetics50-C | VGGSound-C | ||||||||||||
| Gauss. | Traff. | Crowd | Rain | Thund. | Wind | Avg. | Gauss. | Traff. | Crowd | Rain | Thund. | Wind | Avg. | ||
| Source | - | 73.97 | 65.17 | 67.88 | 70.24 | 68.00 | 70.44 | 69.28 | 37.29 | 21.24 | 16.89 | 21.81 | 27.36 | 25.66 | 25.04 |
| TENT* | TTA | 73.02 | 63.36 | 45.31 | 37.02 | 34.57 | 34.01 | 47.88 | 0.68 | 0.28 | 0.28 | 0.28 | 0.28 | 0.28 | 0.35 |
| SAR | TTA | 72.18 | 70.36 | 48.30 | 37.67 | 36.21 | 39.09 | 50.64 | 16.09 | 4.50 | 4.33 | 3.60 | 12.00 | 5.51 | 7.67 |
| CoTTA | CTTA | 19.67 | 4.10 | 2.11 | 2.03 | 2.03 | 2.03 | 5.33 | 5.85 | 1.35 | 0.52 | 0.53 | 0.57 | 0.38 | 1.53 |
| EATA | CTTA | 73.91 | 65.29 | 68.24 | 70.51 | 68.28 | 70.48 | 69.45 | 40.39 | 31.99 | 31.91 | 32.38 | 39.24 | 33.95 | 34.98 |
| MMTTA* | MM-TTA | 17.03 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 4.50 | 0.41 | 0.33 | 0.33 | 0.33 | 0.33 | 0.33 | 0.34 |
| READ | MM-TTA | 68.33 | 59.75 | 57.38 | 54.14 | 53.49 | 52.72 | 57.63 | 18.53 | 7.99 | 7.44 | 5.71 | 8.19 | 4.73 | 8.77 |
| MDAA | MM-CTTA | 72.87 | 71.45 | 72.91 | 72.26 | 73.20 | 73.80 | 72.75 | 38.80 | 34.91 | 34.63 | 34.59 | 37.70 | 35.85 | 36.08 |
| TENT* | TTA | 43.27 | 42.96 | 43.81 | 60.19 | 69.17 | 70.17 | 54.93 | 0.30 | 0.30 | 0.30 | 0.30 | 0.30 | 0.39 | 0.32 |
| SAR | TTA | 41.81 | 27.94 | 24.41 | 40.47 | 42.90 | 70.36 | 41.32 | 14.91 | 4.56 | 4.61 | 3.72 | 12.44 | 5.94 | 7.70 |
| CoTTA | CTTA | 1.99 | 1.99 | 2.92 | 5.07 | 11.64 | 32.56 | 9.36 | 0.30 | 0.30 | 0.30 | 0.30 | 0.30 | 0.36 | 0.31 |
| EATA | CTTA | 73.91 | 65.32 | 68.18 | 70.49 | 68.26 | 70.47 | 69.44 | 40.22 | 33.69 | 31.61 | 32.64 | 39.67 | 32.81 | 35.11 |
| MMTTA* | MM-TTA | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 39.21 | 8.19 | 0.33 | 0.30 | 0.31 | 0.30 | 0.61 | 1.49 | 0.56 |
| READ | MM-TTA | 56.50 | 56.09 | 56.30 | 57.25 | 62.99 | 65.14 | 59.05 | 9.20 | 5.82 | 7.48 | 7.89 | 11.67 | 12.11 | 9.03 |
| MDAA | MM-CTTA | 74.86 | 72.63 | 72.87 | 72.26 | 73.4 | 72.02 | 73.01 | 38.95 | 35.46 | 34.66 | 34.70 | 37.31 | 35.20 | 36.05 |
| Method | Type | Gauss. | Shot. | Impul. | Defoc. | Glass. | Motion. | Zoom. | Snow | Frost | Fog | Bright. | Cont. | Elastic. | Pixel. | Jpeg | Avg. |
| Source | - | 48.74 | 49.80 | 48.99 | 67.68 | 61.84 | 70.88 | 66.18 | 61.35 | 61.39 | 45.34 | 75.95 | 51.87 | 65.77 | 68.78 | 66.10 | 60.71 |
| TENT* | TTA | 16.23 | 2.07 | 2.03 | 2.08 | 2.06 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.98 |
| SAR | TTA | 38.36 | 35.97 | 34.51 | 44.40 | 48.86 | 50.77 | 47.53 | 43.59 | 35.81 | 42.54 | 52.11 | 35.44 | 50.20 | 40.15 | 50.73 | 43.40 |
| CoTTA | CTTA | 33.43 | 27.51 | 25.20 | 21.19 | 18.19 | 16.41 | 14.91 | 13.29 | 11.18 | 9.60 | 8.43 | 6.89 | 6.36 | 5.39 | 4.09 | 14.80 |
| EATA | CTTA | 48.80 | 49.82 | 49.03 | 67.66 | 61.98 | 70.84 | 66.16 | 61.64 | 61.54 | 45.40 | 75.99 | 51.95 | 65.88 | 68.71 | 66.08 | 60.77 |
| MMTTA* | MM-TTA | 14.31 | 2.64 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.89 |
| READ | MM-TTA | 11.92 | 2.04 | 2.03 | 2.97 | 2.41 | 2.46 | 2.41 | 2.30 | 2.04 | 2.04 | 2.04 | 2.04 | 2.04 | 2.04 | 2.03 | 2.86 |
| MDAA | MM-CTTA | 54.89 | 55.25 | 55.32 | 63.89 | 62.49 | 67.26 | 65.86 | 64.32 | 65.31 | 61.86 | 73.20 | 61.60 | 67.83 | 69.22 | 68.69 | 63.80 |
| TENT* | TTA | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.10 | 3.47 | 53.26 | 5.55 |
| SAR | TTA | 34.75 | 35.08 | 35.89 | 42.70 | 45.99 | 49.43 | 50.12 | 44.08 | 42.42 | 40.02 | 57.54 | 35.56 | 48.86 | 57.22 | 66.26 | 45.73 |
| CoTTA | CTTA | 3.89 | 4.01 | 4.50 | 5.39 | 5.47 | 5.75 | 7.82 | 4.98 | 5.79 | 9.85 | 6.89 | 12.72 | 14.26 | 22.33 | 51.34 | 11.00 |
| EATA | CTTA | 48.81 | 49.79 | 49.02 | 67.71 | 61.96 | 70.88 | 66.17 | 61.56 | 61.51 | 45.38 | 75.96 | 51.90 | 65.90 | 68.76 | 66.09 | 60.76 |
| MMTTA* | MM-TTA | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 23.88 | 3.45 |
| READ | MM-TTA | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.19 | 2.93 | 22.95 | 3.50 |
| MDAA | MM-CTTA | 67.32 | 67.48 | 67.76 | 68.98 | 67.60 | 69.59 | 68.49 | 66.46 | 66.18 | 63.14 | 72.87 | 59.33 | 66.59 | 67.64 | 65.25 | 66.98 |
The comparison results of interleaved modality corruption tasks are shown in Table 4 and 5. In this task, EATA, which is more good at memorisation, is not dominant in the task of highlighting reliability bias. READ, which is specifically designed to address intra-modal reliability bias, demonstrates strong performance in this area. However, its effectiveness is limited to video corruption in MM-CTTA, as performance drops significantly during audio corruption. In contrast, MDAA is well-adapted to the corruption of different modalities, outperforming READ by 2.39%-6.84% and EATA by 0.60%-7.28% on average across Kinetics50-C and VGGSound-C.
| Method | Type | Gauss. | Shot. | Impul. | Defoc. | Glass. | Motion. | Zoom. | Snow | Frost | Fog | Bright. | Cont. | Elastic. | Pixel. | Jpeg | Avg. |
| Source | - | 53.02 | 52.90 | 52.98 | 57.20 | 57.38 | 58.37 | 57.48 | 56.40 | 56.46 | 55.41 | 59.16 | 53.73 | 57.22 | 56.44 | 57.33 | 56.10 |
| TENT* | TTA | 51.48 | 50.70 | 50.87 | 51.15 | 50.90 | 51.09 | 50.82 | 50.65 | 50.75 | 50.73 | 50.73 | 50.58 | 50.70 | 50.73 | 50.70 | 50.84 |
| SAR | TTA | 43.12 | 38.99 | 37.77 | 42.43 | 43.84 | 43.61 | 43.79 | 42.13 | 41.26 | 42.83 | 43.84 | 39.34 | 42.75 | 43.49 | 40.38 | 41.97 |
| CoTTA | CTTA | 31.20 | 6.93 | 0.51 | 0.38 | 1.26 | 0.95 | 0.45 | 0.33 | 0.33 | 0.33 | 0.27 | 0.32 | 0.34 | 0.36 | 0.40 | 2.96 |
| EATA | CTTA | 53.57 | 53.70 | 53.57 | 57.00 | 57.29 | 58.46 | 57.77 | 56.24 | 56.57 | 55.45 | 59.06 | 54.13 | 58.24 | 57.22 | 57.38 | 56.38 |
| MMTTA* | MM-TTA | 0.46 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.35 |
| READ | MM-TTA | 33.02 | 0.33 | 0.33 | 0.41 | 0.33 | 0.33 | 0.33 | 0.33 | 0.33 | 0.33 | 0.33 | 0.33 | 0.33 | 0.33 | 0.33 | 2.51 |
| MDAA | MM-CTTA | 55.13 | 55.29 | 55.30 | 56.91 | 57.20 | 57.78 | 57.32 | 56.52 | 56.25 | 56.14 | 58.11 | 55.32 | 57.06 | 56.27 | 57.39 | 56.53 |
| TENT* | TTA | 52.68 | 52.74 | 52.49 | 53.45 | 53.45 | 53.50 | 53.10 | 53.35 | 53.67 | 52.83 | 55.86 | 51.82 | 56.81 | 57.46 | 57.43 | 54.04 |
| SAR | TTA | 40.31 | 39.24 | 38.33 | 41.53 | 41.36 | 44.43 | 43.59 | 42.46 | 41.11 | 41.52 | 42.24 | 38.97 | 43.49 | 40.97 | 46.22 | 41.72 |
| CoTTA | CTTA | 0.33 | 0.33 | 0.35 | 0.33 | 0.33 | 0.33 | 0.32 | 0.46 | 0.36 | 0.35 | 0.56 | 2.24 | 8.36 | 19.22 | 38.95 | 4.85 |
| EATA | CTTA | 53.63 | 53.60 | 53.61 | 57.06 | 57.27 | 58.35 | 57.83 | 56.22 | 56.74 | 55.73 | 59.16 | 54.09 | 58.19 | 57.27 | 57.35 | 56.41 |
| MMTTA* | MM-TTA | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 8.32 | 0.87 |
| READ | MM-TTA | 0.32 | 0.32 | 0.32 | 0.32 | 0.32 | 0.32 | 0.32 | 0.32 | 0.89 | 16.32 | 25.48 | 26.95 | 32.93 | 40.25 | 50.31 | 13.05 |
| MDAA | MM-CTTA | 55.30 | 55.38 | 55.25 | 56.90 | 57.19 | 57.79 | 57.32 | 56.50 | 56.31 | 56.22 | 58.13 | 55.28 | 57.07 | 56.30 | 57.47 | 56.56 |
| Method | V-C | A-C | V-C | A-C | V-C | A-C | V-C | A-C | V-C | A-C | V-C | A-C | V-C | Avg. | ||||||||
| Gauss. | Shot. | Gauss. | Impul. | Defoc. | Traff. | Glass. | Motion. | Crowd | Zoom. | Snow | Frost | Rain | Fog | Bright. | Thund. | Cont. | Elastic. | Wind | Pixel. | Jpeg | ||
| Source | 48.71 | 49.98 | 74.03 | 48.98 | 67.69 | 67.89 | 61.82 | 70.92 | 70.29 | 66.14 | 61.36 | 61.35 | 68.02 | 45.34 | 75.94 | 65.20 | 51.82 | 65.84 | 70.38 | 68.73 | 66.11 | 63.17 |
| TENT* | 48.77 | 48.34 | 74.11 | 46.38 | 62.83 | 67.22 | 62.30 | 68.35 | 69.32 | 64.61 | 54.64 | 57.36 | 66.32 | 46.22 | 63.95 | 37.12 | 38.87 | 42.77 | 40.62 | 10.13 | 5.40 | 51.22 |
| SAR | 48.65 | 49.81 | 74.15 | 48.53 | 66.87 | 65.68 | 62.56 | 70.67 | 68.00 | 66.45 | 58.80 | 60.42 | 69.96 | 47.69 | 75.19 | 67.89 | 50.96 | 65.54 | 70.03 | 66.77 | 63.67 | 62.78 |
| CoTTA | 50.21 | 47.72 | 72.16 | 44.96 | 58.80 | 55.53 | 58.28 | 61.35 | 61.52 | 59.73 | 50.15 | 53.92 | 63.15 | 49.48 | 65.16 | 56.71 | 43.97 | 55.47 | 49.68 | 54.79 | 60.14 | 55.85 |
| EATA | 48.81 | 49.70 | 74.07 | 48.96 | 67.75 | 65.35 | 61.96 | 70.95 | 67.98 | 66.05 | 61.60 | 61.49 | 70.40 | 45.29 | 76.11 | 68.13 | 51.85 | 65.98 | 70.51 | 68.80 | 66.15 | 63.23 |
| MMTTA* | 48.63 | 49.20 | 56.24 | 47.79 | 47.52 | 4.76 | 42.96 | 29.87 | 1.96 | 4.13 | 2.27 | 1.96 | 1.96 | 1.96 | 1.92 | 2.03 | 1.96 | 1.92 | 1.96 | 1.92 | 1.99 | 16.90 |
| READ | 51.18 | 53.62 | 73.88 | 54.73 | 68.67 | 67.95 | 67.07 | 70.14 | 68.84 | 67.62 | 62.68 | 64.90 | 68.24 | 59.46 | 71.43 | 68.16 | 52.60 | 66.35 | 66.18 | 62.85 | 63.35 | 64.28 |
| MDAA | 55.04 | 55.91 | 73.64 | 55.89 | 63.78 | 73.12 | 63.54 | 67.62 | 74.90 | 67.00 | 65.60 | 67.21 | 75.44 | 64.89 | 72.69 | 76.94 | 65.45 | 69.21 | 76.95 | 70.94 | 71.16 | 67.95 |
| Source | 48.73 | 49.75 | 74.02 | 48.99 | 67.61 | 65.21 | 61.93 | 70.87 | 67.87 | 66.17 | 61.36 | 61.43 | 70.27 | 45.32 | 75.88 | 67.97 | 51.84 | 65.74 | 70.47 | 68.74 | 66.10 | 63.16 |
| TENT* | 6.72 | 10.03 | 52.92 | 23.89 | 53.44 | 63.96 | 60.40 | 64.98 | 67.75 | 62.93 | 60.28 | 61.20 | 67.26 | 52.50 | 71.88 | 58.44 | 51.24 | 67.56 | 70.84 | 68.87 | 66.41 | 55.40 |
| SAR | 47.97 | 48.73 | 72.21 | 48.46 | 66.39 | 66.75 | 63.69 | 70.28 | 68.00 | 66.33 | 59.24 | 59.85 | 70.17 | 46.42 | 75.52 | 67.59 | 51.12 | 65.23 | 70.24 | 68.31 | 66.18 | 62.79 |
| CoTTA | 56.40 | 58.83 | 71.10 | 58.56 | 65.11 | 64.63 | 65.33 | 67.54 | 64.49 | 65.16 | 62.44 | 61.61 | 68.98 | 52.40 | 73.39 | 68.37 | 49.00 | 66.00 | 67.99 | 66.45 | 65.62 | 63.78 |
| EATA | 48.76 | 49.76 | 73.89 | 48.86 | 67.75 | 65.20 | 61.89 | 70.67 | 68.28 | 66.16 | 61.64 | 61.44 | 70.33 | 45.57 | 75.96 | 68.30 | 51.88 | 65.87 | 70.51 | 68.59 | 65.98 | 63.20 |
| MMTTA* | 1.96 | 1.96 | 2.00 | 1.96 | 1.96 | 2.00 | 2.00 | 2.03 | 1.96 | 2.03 | 1.96 | 2.00 | 1.96 | 2.52 | 18.31 | 4.88 | 43.11 | 49.71 | 62.28 | 55.69 | 60.04 | 15.35 |
| READ | 51.49 | 52.29 | 71.09 | 50.70 | 62.44 | 63.72 | 65.01 | 66.42 | 66.39 | 64.86 | 60.97 | 64.44 | 67.85 | 62.48 | 74.68 | 72.12 | 54.12 | 69.15 | 70.27 | 69.74 | 68.37 | 64.22 |
| MDAA | 70.44 | 70.21 | 77.35 | 70.12 | 72.13 | 76.56 | 70.05 | 72.44 | 76.07 | 71.46 | 68.94 | 68.72 | 75.92 | 66.51 | 73.32 | 75.10 | 61.78 | 68.05 | 73.52 | 68.56 | 64.91 | 71.06 |
| Method | V-C | A-C | V-C | A-C | V-C | A-C | V-C | A-C | V-C | A-C | V-C | A-C | V-C | Avg. | ||||||||
| Gauss. | Shot. | Gauss. | Impul. | Defoc. | Traff. | Glass. | Motion. | Crowd | Zoom. | Snow | Frost | Rain | Fog | Bright. | Thund. | Cont. | Elastic. | Wind | Pixel. | Jpeg | ||
| Source | 53.05 | 52.91 | 37.32 | 52.98 | 57.19 | 21.24 | 57.37 | 58.37 | 16.89 | 57.45 | 56.37 | 56.47 | 21.82 | 55.41 | 59.19 | 27.37 | 53.75 | 57.19 | 25.66 | 56.44 | 57.33 | 47.23 |
| TENT* | 53.19 | 52.80 | 3.43 | 50.52 | 53.15 | 0.65 | 51.83 | 53.10 | 0.60 | 52.64 | 50.91 | 51.89 | 0.67 | 51.15 | 51.73 | 2.13 | 48.68 | 50.76 | 0.79 | 50.38 | 50.22 | 37.20 |
| SAR | 53.16 | 53.33 | 34.26 | 53.17 | 56.94 | 11.27 | 57.14 | 58.29 | 9.30 | 57.65 | 56.11 | 56.78 | 13.36 | 55.94 | 57.87 | 17.95 | 53.13 | 56.51 | 20.46 | 55.30 | 55.79 | 44.94 |
| CoTTA | 52.24 | 52.26 | 10.66 | 49.15 | 49.22 | 1.68 | 46.15 | 46.38 | 1.34 | 44.84 | 44.66 | 44.13 | 0.65 | 43.57 | 43.23 | 4.83 | 35.16 | 34.85 | 0.56 | 36.10 | 35.34 | 32.24 |
| EATA | 53.63 | 53.70 | 40.51 | 53.59 | 57.21 | 30.81 | 57.45 | 58.49 | 29.80 | 57.85 | 56.37 | 56.85 | 30.55 | 56.72 | 59.13 | 37.29 | 54.31 | 58.27 | 32.58 | 57.28 | 57.55 | 50.00 |
| MMTTA* | 8.27 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.72 |
| READ | 53.78 | 53.91 | 39.83 | 54.17 | 57.81 | 26.00 | 58.14 | 59.42 | 21.63 | 59.03 | 57.38 | 58.29 | 22.79 | 57.71 | 59.32 | 26.07 | 55.46 | 58.36 | 18.29 | 57.36 | 57.63 | 48.21 |
| MDAA | 55.09 | 55.31 | 38.60 | 55.31 | 56.89 | 34.83 | 57.20 | 57.69 | 34.65 | 57.35 | 56.47 | 56.27 | 34.28 | 56.17 | 58.05 | 36.86 | 55.33 | 57.00 | 35.53 | 56.28 | 57.35 | 50.60 |
| Source | 53.01 | 52.88 | 37.31 | 52.97 | 57.20 | 21.25 | 57.42 | 58.41 | 16.89 | 57.49 | 56.37 | 56.49 | 21.81 | 55.43 | 59.16 | 27.37 | 53.74 | 57.19 | 25.66 | 56.42 | 57.29 | 47.23 |
| TENT* | 24.80 | 42.44 | 1.65 | 49.78 | 51.65 | 0.32 | 51.64 | 51.52 | 0.34 | 51.35 | 50.51 | 50.89 | 0.43 | 51.27 | 52.07 | 1.14 | 51.14 | 54.94 | 1.64 | 56.50 | 56.92 | 35.85 |
| SAR | 51.90 | 51.78 | 25.90 | 51.30 | 54.73 | 5.11 | 54.82 | 56.52 | 7.82 | 56.31 | 54.63 | 55.49 | 13.44 | 55.12 | 57.59 | 15.07 | 53.57 | 56.92 | 15.04 | 56.53 | 57.30 | 43.19 |
| CoTTA | 42.97 | 43.65 | 6.10 | 44.29 | 44.97 | 1.09 | 45.45 | 45.60 | 2.20 | 49.53 | 48.97 | 49.89 | 13.53 | 51.33 | 53.72 | 19.60 | 52.45 | 56.04 | 18.76 | 56.22 | 55.88 | 38.20 |
| EATA | 53.77 | 53.65 | 40.39 | 53.61 | 57.17 | 30.49 | 57.36 | 58.59 | 30.16 | 57.82 | 56.21 | 56.64 | 31.22 | 56.77 | 59.24 | 37.44 | 54.30 | 58.18 | 33.01 | 57.46 | 57.53 | 50.05 |
| MMTTA* | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 8.36 | 18.74 | 46.32 | 3.79 |
| READ | 54.35 | 54.56 | 25.03 | 54.38 | 57.99 | 17.67 | 57.76 | 58.74 | 20.57 | 58.66 | 56.91 | 57.58 | 20.81 | 58.04 | 59.10 | 33.82 | 55.54 | 58.13 | 32.75 | 57.80 | 58.34 | 48.03 |
| MDAA | 55.30 | 55.41 | 38.64 | 55.29 | 56.91 | 35.40 | 57.14 | 57.81 | 34.85 | 57.34 | 56.48 | 56.29 | 34.52 | 56.23 | 58.24 | 37.22 | 55.31 | 57.13 | 35.12 | 56.30 | 57.47 | 50.69 |
4.3 Ablation Study
In this section, we conduct three ablation studies on both video-corrupted Kinetics50-C dataset and audio-corrupted VGGSound-C dataset in severity level 5, with the batch size of 64. For simplicity, in the following section we use KS-video and VGG-audio to represent these two tasks.
Component analysis. To verify the effectiveness of each MDAA component, we adopt an ablation study w.r.t three components as shown in Table 6. As observed, the model using only AC underperforms, with an average accuracy 0.5%-0.57% lower than READ on KS-video, and the model even collapses on VGG-audio, with an average accuracy of 0.47%-0.56%. This occurs because, while AC can prevent model forgetting, it cannot filter out unreliable samples, leading to issues with error accumulation and reliability bias. In contrast, the addition of DSM significantly improves performance, with gains of 1.73%-7.52% on KS-video and 33.74%-33.90% on VGG-audio. The introduction of SPS allows the model to learn from more possibly correct labels at the same time, thus further improving the performances on most tasks.
| Method | KS-video | VGG-audio | ||
| READ | 62.32 | 62.59 | 23.93 | 22.39 |
| MDAA (ACs) | 61.82 | 62.02 | 0.47 | 0.56 |
| MDAA (ACs+DSM) | 63.55 | 69.54 | 34.87 | 34.85 |
| MDAA (ACs+DSM+SPS) | 65.43 | 69.30 | 35.82 | 35.77 |
Reliable selection threshold. To examine the effect of the threshold on the DSM, we plot the model’s performance with of 0, 1-4, 5-4, 1-3, 2-3 and 5-3 in Fig.4(A-B). The performance of the model on both datasets exhibits an increasing and then decreasing trend. When is close to , the ACs are updated for nearly every sample, introducing more error. Conversely, when increases too much, the ACs do not learn from new inputs since no samples can pass through the DSM, leading to a decline in performance. Therefore, the we choose 1-3 a moderate valuefor both datasets as the threshold.
Soft label reconstruction. In this part we determine a suitable number to reconstruct the pseudo-labels. Given the sorted top distribution set ,we assign weights to each class in a decreasing manner using the formula . The results of the model are plotted in Fig.4(C-D) with chosen from to . While using soft labels can inevitably introduce error, there are instances where performance with one-hot labels may exceed that of the SPS (e.g., KS-video ). However, SPS remains beneficial as the dataset becomes larger and more complex. In fact, performance using SPS surpasses that of one-hot labels in most tasks, making SPS a worthwhile trade-off in MM-CTTA. Generally, the optimal number for reconstructions depends on the number of categories in the dataset, and a larger number of categories warrants a larger . However, more classes will be included in the reconstruction as increase, thereby introducing more error and reducing model performance. Therefore in SPS, we determine to use top and classes to reconstructed label in Kinetics50-C and VGGSound-C respectively.
5 Conclusion
In this paper, we analysed the factors that affect the model in the MM-CTTA task (i.e., error accumulation, catastrophic forgetting and reliability bias) and demonstrate that typical TTA methods are not suitable for the MM-CTTA task. To address the impact of these factors, we propose a new paradigm MDAA that introduce analytic learning to TTA for the first time. Instead of just adapting the model to the target domain, MDAA integrates the target domain into source domain, and thus prevent network from forgetting. With the help of DSM and SPS, model is able to dynamically and comprehensively process the information provided by each modality and use reliable samples to update. In the future, we will try to adapt this paradigm to more modalities to solve more challenging problems in real scenarios.
Appendix A Proofs of Theorems
In this section, we provide comprehensive proofs of Theorems 1 and 2 presented in the main paper.
Note above equation as , derive for as
Therefore .
Proofs of Theorem 2. Similar to the proof of Theorem 1, we note the optimal formula as and derive it in terms of as
Therefore .
Appendix B Pseudo-code for MDAA
In this appendix, we provide the pseudo-code for our MDAA in Algorithm 1. For the pre-trained model, we integrate an individual Analytic Classifier (AC) for each network block, using the source dataset to initialize the classifiers as well as the memory bank. During the inference and adaptation periods, the model reconstructs the output labels for each sample using the Soft Pseudo-label Strategy (SPS) and determines which ACs need to be updated through the Dynamic Selecting Mechanism (DSM).
Appendix C Benchmarks,Backbone and Implementation Details
C.1 Details about the Benchmarks
ALL experiments are conducted on the two popular multi-modal datasets Kinetics (Kay et al., 2017) and VGGSound (Chen et al., 2020). Yang et al. (2024) further provides their corrupted visual and audio modality for TTA tasks.
Kinetics50 is a subset of the Kinetics dataset (Kay et al., 2017), consisting of 50 randomly selected classes (Yang et al., 2024). It primarily includes videos that focus on human motion-related activities, with each clip lasting approximately 10 seconds and labeled with a single action class. All videos are sourced from YouTube. The Kinetics50 dataset comprises 29,204 visual-audio pairs for training and 2,466 pairs for testing, with the video modality playing a more prominent role in modality pairing.
VGGSound is a large-scale audio-visual dataset containing short audio clips extracted from YouTube videos (Chen et al., 2020), covering 309 distinct everyday audio events. Each clip has a fixed duration of 10 seconds. For this study, we utilize the 157,602 pairs for training and 14,046 pairs for testing. Compared to Kinetics50, VGGSound includes a wider range of categories, introducing additional complexity to the classification task.
Both datasets’ visual and audio modalities were extracted following the method described in Gong et al. (2022b). To systematically explore the distributional shifts across modalities, various corruption types were applied to both visual and audio components. Following Yang et al. (2024), 15 corruption types were introduced for the visual modality, each with five levels of severity for comprehensive evaluation. These corruptions include ”gaussian noise”, ”shot noise”, ”impulse noise”, ”defocus blur”, ”glass blur”, ”motion blur”, ”zoom blur”, ”snow”, ”frost”, ”fog”, ”brightness”, ”contrast”, ”elastic transform”, ”pixelate”, and ”jpeg compression”. Similarly, the audio modality was subjected to 6 different corruptions: ”gaussian noise”, ”traffic noise”, ”crowd noise”, ”rain”, ”thunder”, and ”wind”. The corrupted versions of these benchmarks are referred to as Kinetics50-C and VGGSound-C, respectively. Visualizations of sample corrupted video frames and audio spectrograms are provided in Fig.5 and Fig.6.
C.2 CAV-MAE Backbone
CAV-MAE is employed as the pre-trained model for MM-CTTA in this paper. Its architecture consists of 11 Transformer blocks (known as feature encoder networks) dedicated to modality-specific feature extraction, followed by an additional Transformer block (known as fusion network) responsible for cross modal fusion. For the video input, 10 frames are sampled from each clip, from which a single frame is randomly selected and fed into the Transformer encoder for the visual modality. In the case of the audio input, the original 10-second audio waveform is transformed into a 2D spectrogram before being processed by the Transformer encoder for the audio modality (Gong et al., 2022a).
C.3 Implementation Details
In the final step of Algorithm 1, we determine the hyperparameters for MDAA. The expansion layer dimension, denoted as , theoretically benefits from larger values. However, an excessively large dimension may introduce a significant number of parameters, increasing the computational load. Given the constraints of our available GPU resources, we set the dimension of to 8000. The necessity of the parameter in 3.2 has been established in Zhuang et al. (2022). The model demonstrates stable performance over a wide range of values, indicating that as long as is within a reasonable range, its impact on the model’s performance remains minimal. After conducting a sweep over five orders of magnitude (), we set to 1 for Kinetics-50 and 10 for VGGSound. As discussed in Sections 4.2 and 4.3, the threshold in DSM is fixed at 0.001 for both datasets, while the parameter in SPS is set to 2 for Kinetics-50 and 7 for VGGSound. All experiments were conducted on an RTX3090 GPU, with results averaged over three runs.
Appendix D More Experiment Results
In this appendix we provide four more experiments for reference where Appendix .D.1 and D.2 are the supplementary comparative studies while Appendix .D.3 and D.4 are the further discussion on hyperparameters applied in MDAA.
D.1 Performance Comparison on Multi-modal Corruption
We follow the experimental setup in (Yang et al., 2024) to examine whether our model remains superior to others when the adaptation task is not continual. In this section, the corrupted target domains are treated as independent, and the results for corrupted audio and video modalities are presented in Tables 7, 8, and 9. The results show that the SOTA performance of MDAA is not only due to its ability to combat catastrophic forgetting but also its strong capacity to handle MM-TTA tasks effectively.
| Method | Type | Kinetics50-C | VGGSound-C | ||||||||||||
| Gauss. | Traff. | Crowd | Rain | Thund. | Wind | Avg. | Gauss. | Traff. | Crowd | Rain | Thund. | Wind | Avg. | ||
| Source | - | 73.7 | 65.5 | 67.9 | 70.3 | 67.9 | 70.3 | 69.3 | 37.0 | 25.5 | 16.8 | 21.6 | 27.3 | 25.5 | 25.6 |
| TENT | TTA | 73.9 | 67.4 | 69.2 | 70.4 | 66.5 | 70.5 | 69.6 | 10.6 | 2.6 | 1.8 | 2.8 | 5.3 | 4.1 | 4.5 |
| SAR | TTA | 73.7 | 65.4 | 68.2 | 69.9 | 67.2 | 70.2 | 69.1 | 37.4 | 9.5 | 11.0 | 12.1 | 26.8 | 23.7 | 20.1 |
| EATA | CTTA | 73.7 | 66.1 | 68.5 | 70.3 | 67.9 | 70.1 | 69.4 | 39.2 | 26.1 | 22.9 | 26.0 | 31.7 | 30.4 | 29.4 |
| MMTTA | MM-TTA | 70.8 | 69.2 | 68.5 | 69.0 | 69.8 | 69.4 | 69.4 | 14.1 | 5.2 | 6.4 | 6.9 | 8.6 | 4.5 | 7.6 |
| READ | MM-TTA | 74.1 | 69.0 | 69.7 | 71.1 | 71.8 | 70.7 | 71.1 | 40.4 | 28.9 | 26.6 | 30.9 | 36.7 | 30.6 | 32.4 |
| MDAA | MM-CTTA | 73.8 | 70.3 | 71.0 | 70.9 | 72.8 | 71.4 | 71.7 | 38.6 | 34.9 | 34.6 | 34.3 | 37.4 | 35.2 | 35.8 |
| Method | Type | Gauss. | Shot. | Impul. | Defoc. | Glass. | Motion. | Zoom. | Snow | Frost | Fog | Bright. | Cont. | Elastic. | Pixel. | Jpeg | Avg. |
| Source | - | 46.8 | 48.0 | 46.9 | 67.5 | 62.2 | 70.8 | 66.7 | 61.6 | 60.3 | 46.7 | 75.2 | 52.1 | 65.7 | 66.5 | 61.9 | 59.9 |
| TENT | TTA | 46.3 | 47.0 | 46.3 | 67.2 | 62.5 | 71.0 | 67.6 | 63.1 | 61.1 | 34.9 | 75.4 | 51.6 | 66.8 | 67.2 | 62.7 | 59.4 |
| SAR | TTA | 46.7 | 47.4 | 46.8 | 67.0 | 61.9 | 70.4 | 66.4 | 61.8 | 60.6 | 46.0 | 75.2 | 52.1 | 65.7 | 66.4 | 62.0 | 59.8 |
| EATA | CTTA | 46.8 | 47.6 | 47.1 | 67.2 | 62.7 | 70.6 | 67.2 | 62.3 | 60.9 | 46.7 | 75.2 | 52.4 | 65.9 | 66.8 | 62.5 | 60.1 |
| MMTTA | MM-TTA | 46.2 | 46.6 | 46.1 | 58.8 | 55.7 | 62.6 | 58.7 | 52.6 | 54.4 | 48.5 | 69.1 | 49.3 | 57.6 | 56.4 | 54.6 | 54.5 |
| READ | MM-TTA | 49.4 | 49.7 | 49.0 | 68.0 | 65.1 | 71.2 | 69.0 | 64.5 | 64.4 | 57.4 | 75.5 | 53.6 | 68.3 | 68.0 | 65.1 | 62.5 |
| MDAA | MM-CTTA | 55.1 | 55.3 | 55.7 | 64.5 | 62.3 | 67.7 | 65.0 | 61.6 | 63.6 | 57.9 | 72.2 | 54.8 | 66.6 | 67.0 | 65.2 | 62.3 |
| Method | Type | Gauss. | Shot. | Impul. | Defoc. | Glass. | Motion. | Zoom. | Snow | Frost | Fog | Bright. | Cont. | Elastic. | Pixel. | Jpeg | Avg. |
| Source | - | 52.8 | 52.7 | 52.7 | 57.2 | 57.2 | 58.7 | 57.6 | 56.4 | 56.6 | 55.6 | 58.9 | 53.7 | 56.9 | 55.8 | 56.9 | 56.0 |
| TENT | TTA | 52.7 | 52.7 | 52.7 | 56.7 | 56.5 | 57.9 | 57.2 | 55.9 | 56.3 | 56.3 | 58.4 | 54.0 | 57.4 | 56.2 | 56.7 | 55.8 |
| SAR | TTA | 52.9 | 52.8 | 52.9 | 57.2 | 57.1 | 58.6 | 57.6 | 56.3 | 56.7 | 55.9 | 58.9 | 54.0 | 57.0 | 56.0 | 57.0 | 56.1 |
| EATA | CTTA | 53.0 | 52.8 | 53.0 | 57.2 | 57.1 | 58.6 | 57.8 | 56.3 | 56.8 | 56.4 | 59.0 | 54.1 | 57.4 | 56.1 | 57.0 | 56.2 |
| MMTTA | MM-TTA | 7.1 | 7.3 | 7.3 | 44.8 | 41.5 | 48.0 | 45.5 | 27.4 | 23.5 | 30.5 | 46.9 | 24.2 | 40.3 | 40.7 | 45.7 | 32.0 |
| READ | MM-TTA | 53.6 | 53.6 | 53.5 | 57.9 | 57.7 | 59.4 | 58.8 | 57.2 | 57.8 | 55.0 | 59.9 | 55.2 | 58.6 | 57.1 | 57.9 | 56.9 |
| MDAA | MM-CTTA | 54.89 | 55.25 | 55.32 | 63.89 | 62.49 | 67.26 | 65.86 | 64.32 | 65.31 | 61.86 | 73.20 | 61.60 | 67.83 | 69.22 | 68.69 | 63.80 |
D.2 Performance Comparison on Single-modality Continual Corruption
In this section, we compare the performance of each method under single-modality corruption. This task is similar to the progressive single-modality corruption task described in the main text, but here we use a batch size of 64.
| Method | Type | Kinetics50-C | VGGSound-C | ||||||||||||
| Gauss. | Traff. | Crowd | Rain | Thund. | Wind | Avg. | Gauss. | Traff. | Crowd | Rain | Thund. | Wind | Avg. | ||
| Source | - | 73.97 | 65.21 | 67.79 | 70.27 | 67.98 | 70.45 | 69.28 | 37.32 | 21.24 | 16.89 | 21.82 | 27.37 | 25.66 | 25.05 |
| TENT* | TTA | 74.44 | 68.04 | 71.30 | 70.25 | 72.53 | 70.35 | 71.15 | 10.76 | 1.15 | 0.40 | 0.32 | 0.51 | 0.31 | 2.24 |
| SAR | TTA | 73.88 | 65.68 | 68.00 | 70.91 | 69.07 | 70.45 | 69.66 | 37.39 | 8.57 | 7.03 | 12.58 | 10.77 | 13.71 | 15.01 |
| CoTTA | CTTA | 73.89 | 66.84 | 68.08 | 67.53 | 71.10 | 69.33 | 69.46 | 39.70 | 34.88 | 35.54 | 33.67 | 39.42 | 33.83 | 36.17 |
| EATA | CTTA | 73.95 | 65.26 | 68.03 | 70.45 | 68.17 | 70.48 | 69.39 | 40.49 | 31.07 | 31.98 | 31.40 | 38.26 | 33.84 | 34.51 |
| MMTTA* | MM-TTA | 69.32 | 69.01 | 69.00 | 69.07 | 68.96 | 68.95 | 69.05 | 14.40 | 1.92 | 0.84 | 0.36 | 0.47 | 0.31 | 3.05 |
| READ | MM-TTA | 74.74 | 68.88 | 70.43 | 70.69 | 72.31 | 69.73 | 71.13 | 40.51 | 25.39 | 20.38 | 20.06 | 21.14 | 16.07 | 23.93 |
| MDAA | MM-CTTA | 73.33 | 71.99 | 73.36 | 73.26 | 74.24 | 73.76 | 73.32 | 38.57 | 34.57 | 34.37 | 34.30 | 37.40 | 35.70 | 35.82 |
| Source | - | 74.00 | 65.20 | 67.92 | 70.24 | 68.01 | 70.43 | 69.30 | 37.31 | 21.25 | 16.89 | 21.81 | 27.37 | 25.66 | 25.05 |
| TENT* | TTA | 73.31 | 70.00 | 71.57 | 70.30 | 68.87 | 71.09 | 70.86 | 0.37 | 0.30 | 0.30 | 0.31 | 0.97 | 3.73 | 1.00 |
| SAR | TTA | 73.29 | 66.79 | 68.14 | 70.71 | 67.97 | 70.36 | 69.55 | 32.02 | 9.92 | 7.04 | 9.49 | 11.41 | 16.03 | 14.32 |
| CoTTA | CTTA | 69.88 | 67.67 | 67.84 | 68.54 | 69.63 | 70.76 | 69.05 | 25.06 | 25.08 | 28.01 | 33.22 | 37.90 | 29.67 | 29.82 |
| EATA | CTTA | 73.96 | 65.30 | 68.15 | 70.36 | 68.18 | 70.40 | 69.39 | 40.65 | 32.32 | 30.65 | 32.23 | 38.16 | 32.39 | 34.40 |
| MMTTA* | MM-TTA | 68.92 | 69.47 | 70.50 | 69.59 | 69.62 | 69.63 | 69.62 | 0.35 | 0.32 | 0.30 | 0.23 | 0.67 | 3.25 | 0.85 |
| READ | MM-TTA | 73.11 | 70.16 | 69.68 | 70.90 | 72.07 | 70.73 | 71.11 | 24.23 | 16.23 | 16.84 | 17.65 | 29.12 | 30.30 | 22.39 |
| MDAA | MM-CTTA | 72.55 | 73.67 | 72.98 | 73.04 | 73.18 | 71.55 | 72.83 | 38.59 | 35.35 | 34.50 | 34.31 | 36.94 | 34.91 | 35.77 |
| Method | Type | Gauss. | Shot. | Impul. | Defoc. | Glass. | Motion. | Zoom. | Snow | Frost | Fog | Bright. | Cont. | Elastic. | Pixel. | Jpeg | Avg. |
| Source | - | 48.67 | 49.81 | 49.01 | 67.77 | 61.88 | 70.95 | 66.19 | 61.39 | 61.42 | 45.35 | 75.94 | 51.86 | 65.81 | 68.77 | 66.10 | 60.73 |
| TENT* | TTA | 48.53 | 48.65 | 46.06 | 62.91 | 62.15 | 65.77 | 63.83 | 54.78 | 54.64 | 33.79 | 36.79 | 19.86 | 11.75 | 3.82 | 3.38 | 41.11 |
| SAR | TTA | 48.48 | 49.87 | 48.71 | 66.92 | 62.56 | 70.58 | 66.77 | 59.25 | 60.50 | 47.19 | 75.34 | 50.77 | 65.23 | 66.91 | 64.01 | 60.21 |
| CoTTA | CTTA | 49.17 | 46.65 | 43.69 | 61.82 | 60.00 | 63.54 | 62.34 | 50.60 | 52.13 | 53.35 | 60.25 | 50.86 | 60.43 | 58.28 | 62.61 | 55.71 |
| EATA | CTTA | 48.76 | 49.84 | 49.03 | 67.78 | 62.02 | 70.92 | 66.20 | 61.55 | 61.53 | 45.38 | 75.97 | 51.90 | 65.91 | 68.74 | 66.09 | 60.77 |
| MMTTA* | MM-TTA | 48.74 | 49.05 | 48.88 | 49.12 | 48.94 | 48.88 | 48.86 | 48.92 | 48.92 | 48.86 | 49.03 | 48.79 | 48.88 | 48.80 | 48.92 | 48.91 |
| READ | MM-TTA | 51.02 | 53.54 | 54.24 | 68.16 | 66.36 | 68.95 | 67.76 | 62.85 | 64.72 | 59.62 | 70.99 | 53.02 | 67.03 | 63.71 | 62.78 | 62.32 |
| MDAA | MM-CTTA | 55.84 | 55.79 | 55.50 | 64.29 | 63.71 | 68.04 | 68.10 | 66.06 | 68.01 | 65.71 | 72.97 | 66.36 | 70.04 | 70.51 | 70.53 | 65.43 |
| Source | - | 48.68 | 49.79 | 48.97 | 67.69 | 61.82 | 70.93 | 66.18 | 61.39 | 61.37 | 45.29 | 75.99 | 51.89 | 65.78 | 68.71 | 66.14 | 60.71 |
| TENT* | TTA | 41.93 | 47.19 | 49.22 | 64.80 | 64.33 | 68.05 | 63.86 | 61.42 | 62.12 | 51.65 | 73.92 | 50.80 | 67.47 | 69.03 | 66.80 | 60.17 |
| SAR | TTA | 48.08 | 49.07 | 48.83 | 66.25 | 63.02 | 70.39 | 66.27 | 59.53 | 60.31 | 46.28 | 75.62 | 51.23 | 65.47 | 68.47 | 66.14 | 60.33 |
| CoTTA | CTTA | 56.89 | 57.26 | 56.92 | 66.39 | 65.16 | 68.11 | 66.52 | 60.23 | 62.46 | 50.36 | 73.47 | 50.02 | 66.04 | 67.85 | 67.06 | 62.31 |
| EATA | CTTA | 48.67 | 49.80 | 49.16 | 67.68 | 62.02 | 70.92 | 66.13 | 61.57 | 61.46 | 45.17 | 76.02 | 51.95 | 65.90 | 68.71 | 66.22 | 60.76 |
| MMTTA* | MM-TTA | 49.82 | 49.90 | 49.85 | 49.94 | 49.86 | 49.92 | 49.89 | 49.97 | 49.94 | 49.79 | 50.54 | 50.01 | 54.33 | 56.20 | 59.99 | 51.33 |
| READ | MM-TTA | 52.00 | 52.28 | 51.02 | 62.43 | 64.47 | 66.04 | 65.38 | 61.39 | 65.10 | 62.96 | 75.14 | 53.59 | 68.89 | 69.95 | 68.20 | 62.59 |
| MDAA | MM-CTTA | 70.55 | 70.43 | 70.08 | 71.81 | 70.30 | 72.19 | 70.88 | 69.38 | 69.35 | 67.48 | 73.16 | 62.07 | 68.63 | 67.98 | 65.21 | 69.30 |
| Method | Type | Gauss. | Shot. | Impul. | Defoc. | Glass. | Motion. | Zoom. | Snow | Frost | Fog | Bright. | Cont. | Elastic. | Pixel. | Jpeg | Avg. |
| Source | - | 53.05 | 52.91 | 52.98 | 57.19 | 57.37 | 58.37 | 57.45 | 56.37 | 56.47 | 55.41 | 59.19 | 53.75 | 57.19 | 56.44 | 57.33 | 56.10 |
| TENT* | TTA | 53.27 | 52.76 | 52.00 | 54.58 | 54.35 | 55.04 | 54.86 | 52.59 | 52.81 | 53.11 | 53.50 | 50.80 | 53.03 | 52.31 | 52.14 | 53.14 |
| SAR | TTA | 53.14 | 53.29 | 53.21 | 56.95 | 56.91 | 58.41 | 57.47 | 56.10 | 56.78 | 56.34 | 58.35 | 53.98 | 57.27 | 55.93 | 56.48 | 56.04 |
| CoTTA | CTTA | 53.03 | 53.08 | 53.03 | 52.92 | 52.28 | 51.84 | 51.41 | 50.83 | 50.08 | 49.66 | 49.22 | 48.85 | 48.30 | 47.62 | 47.20 | 50.62 |
| EATA | CTTA | 53.74 | 53.67 | 53.68 | 57.20 | 57.26 | 58.53 | 57.93 | 56.38 | 56.67 | 56.23 | 59.04 | 53.63 | 58.19 | 57.36 | 57.48 | 56.47 |
| READ | MM-TTA | 53.77 | 54.26 | 54.28 | 58.04 | 58.00 | 59.09 | 58.84 | 57.43 | 58.18 | 58.12 | 59.38 | 55.99 | 58.30 | 57.51 | 57.91 | 57.27 |
| MDAA | MM-CTTA | 55.63 | 55.91 | 55.88 | 57.50 | 57.77 | 58.27 | 57.84 | 57.05 | 56.72 | 56.77 | 58.67 | 55.78 | 57.55 | 56.87 | 57.84 | 57.07 |
| Source | - | 53.01 | 52.88 | 52.97 | 57.20 | 57.42 | 58.41 | 57.49 | 56.37 | 56.49 | 55.43 | 59.16 | 53.74 | 57.19 | 56.42 | 57.29 | 56.10 |
| TENT* | TTA | 50.48 | 50.50 | 50.50 | 53.09 | 53.19 | 53.70 | 53.18 | 52.31 | 53.47 | 53.57 | 55.40 | 52.22 | 56.23 | 56.74 | 56.99 | 53.44 |
| SAR | TTA | 53.02 | 52.95 | 52.71 | 56.78 | 56.63 | 58.18 | 57.38 | 55.97 | 56.69 | 56.45 | 58.46 | 54.03 | 57.41 | 56.53 | 57.16 | 56.02 |
| CoTTA | CTTA | 48.10 | 48.51 | 49.13 | 49.90 | 50.47 | 50.93 | 51.67 | 52.14 | 53.17 | 53.19 | 55.86 | 53.21 | 58.02 | 56.60 | 56.71 | 52.51 |
| EATA | CTTA | 53.69 | 53.66 | 53.61 | 57.26 | 57.34 | 58.46 | 57.84 | 56.35 | 56.82 | 56.72 | 59.21 | 54.10 | 58.28 | 57.39 | 57.54 | 56.55 |
| READ | MM-TTA | 55.33 | 55.34 | 54.93 | 58.16 | 57.88 | 58.82 | 58.50 | 57.24 | 57.94 | 57.86 | 59.65 | 55.67 | 58.79 | 57.87 | 58.10 | 57.47 |
| MDAA | MM-CTTA | 55.85 | 55.87 | 55.94 | 57.40 | 57.85 | 58.25 | 57.95 | 57.02 | 56.89 | 56.68 | 58.66 | 55.84 | 57.58 | 56.92 | 58.00 | 57.11 |
D.3 Dynamic threshold update
The threshold we used in DSM is a fixed number. In this section we attempt to update in a dynamic way during the adaptation. We define the threshold for classifier in time as
where is the learning rate, is the initial threshold and is the batch size. is calculated to reflect the gap between leader and follower . The original intention of this design is to adjust the size of the threshold according to the change of , so as to eliminate the statistical bias of different domains. However, as shown in Table 14, such attempt is not only achieve lower performance while needs more variables to be memorized. So only the fixed threshold is used in the formal method.
| KS-video | VGG-audio | |||
| 0 | 63.55 | 69.54 | 34.37 | 34.35 |
| 0.01 | 63.49 | 69.31 | 34.35 | 34.30 |
| 0.05 | 63.36 | 69.43 | 34.38 | 34.31 |
| 0.1 | 63.18 | 69.42 | 34.40 | 34.29 |
| 0.2 | 62.93 | 69.33 | 34.41 | 34.31 |
| KS-video | ||
| 0.9 | 64.29 | 69.29 |
| 0.8 | 64.88 | 69.29 |
| 0.7 | 65.43 | 69.30 |
| 0.6 | 65.20 | 68.81 |
| 0.5 | 35.37 | 37.37 |
D.4 Label weight distribution
The weight assignment in SPS follows the formula . In this section we make a toy experiment to explore whether such assignment is reasonable. We take KS-video as example which use top 2 classes for reconstruction. Table 14 show the results on different weight assignment. It can be seen that the performance peaks at 0.7, corresponding to the assignment in main text. Therefore the assignment of weights in the main text is justified.

### v2

@@
 Analytic Continual Test-Time Adaptation for Multi-Modality Corruption
-Abstract
-Test-Time Adaptation (TTA) aims to help pre-trained model bridge the gap between source and target datasets using only the pre-trained model and unlabelled test data.
-A key objective of TTA is to address domain shifts in test data caused by corruption, such as weather changes, noise, or sensor malfunctions.
-Multi-Modal Continual Test-Time Adaptation (MM-CTTA), an extension of TTA with better real-world applications, further allows pre-trained models to handle multi-modal inputs and adapt to continuously-changing target domains.
-MM-CTTA typically faces challenges including error accumulation, catastrophic forgetting, and reliability bias, with few existing approaches effectively addressing these issues in multi-modal corruption scenarios.
-In this paper, we propose a novel approach, Multi-modality Dynamic Analytic Adapter (MDAA), for MM-CTTA tasks.
-We innovatively introduce analytic learning into TTA, using the Analytic Classifiers (ACs) to prevent model forgetting.
-Additionally, we develop Dynamic Selection Mechanism (DSM) and Soft Pseudo-label Strategy (SPS), which enable MDAA to dynamically filter reliable samples and integrate information from different modalities.
-Extensive experiments demonstrate that MDAA achieves state-of-the-art performance on MM-CTTA tasks while ensuring reliable model adaptation.
-1 Introduction
-Test-Time Adaptation (TTA) aims to help the pre-trained model bridge the gap between the source domain and the target domain (Wang et al., 2021; Liang et al., 2024).
-Unlike Unsupervised Domain Adaptation (UDA) (Zhang et al., 2015; Liang et al., 2024), TTA performs adaptation without the need for any source data (i.e., pre-trained dataset), which not only saves computational resources by avoiding retraining but also preserves the privacy of the source data.
-One key TTA application is addressing the problem of domain shift from source data to corrupted test data, where the corruption is often caused by external factors (e.g., weather changes, ambient noise) or sensor malfunctions.
-As an extension of TTA, Continual Test-Time Adaptation (CTTA) has been proposed to align with real-world scenarios where domain shifts usually are dynamic (Wang et al., 2022).
-Challenges in CTTA mainly consist of error accumulation and catastrophic forgetting.
-Error accumulation, stemming from incorrect pseudo-labels, can mislead models’ adaptation and potentially lead to collapse (Chen et al., 2019).
-Catastrophic forgetting refers to the loss of knowledge from the source data during continuous adaptation, reducing the model’s generalization ability (McCloskey & Cohen, 1989).
-To address these challenges, various CTTA methods have been proposed, yielding promising results in corruption-related tasks (Wang et al., 2022; Gao et al., 2023; Niu et al., 2022).
-Most existing CTTA approaches focus solely on single-modal scenarios, paying less attention to multi-modal applications.
-Compared with TTA and CTTA, Multi-Modal Continual Test-Time Adaptation (MM-CTTA) (Cao et al., 2023) shows greater potential for real-world applications, as multi-modal data integrates a broader range of information than single-modality adaptation, resulting in more robust networks (Radford et al., 2021).
-However, applying existing CTTA methods to MM-TTA by simply replacing the backbone with muiti-modal encoders is less optimal.
-MM-CTTA performance can easily suffer from reliability bias, where intra-modality domain shifts increase information discrepancies in downstream fusion networks (Yang et al., 2024).
-Such effect becomes more pronounced when modality corruption changes dynamically.
-Although a few works address Multi-Modal Test-Time Adaptation (MM-TTA) (Yang et al., 2024; Lei & Pernkopf, 2024), they struggle with model forgetting during continuous adaptation and the challenges posed by interleaved modality corruption.
-To further illustrate, we use a representative TTA method, TENT (Wang et al., 2021), as an example in Fig.1.
-The results indicate that traditional TTA methods face significant limitations due to error accumulation, catastrophic forgetting, and reliability bias.
-These challenges suggest that methods designed for MM-CTTA must meet three key requirements: (1) effectively mitigating error accumulation during adaptation, (2) retaining knowledge of the source data after adapting to various domains, and (3) dynamically suppressing the influence of corrupted modalities while prioritizing more reliable ones.
-To the best of our knowledge, the method closest to meeting these criteria is CoMAC (Cao et al., 2023), which focuses on segmentation tasks in living environments rather than addressing modality corruption.
-In this paper, we propose a new approach named the Multi-modality Dynamic Analytic Adaptor (MDAA) to address the challenges in MM-CTTA.
-MDAA comprises three primary components: (i) the Analytic Classifiers (ACs), (ii) the Dynamic Selection Mechanism (DSM), and (iii) the Soft Pseudo-label Strategy (SPS).
-ACs update the model by addressing a recursive ridge regression problem, optimizing on both new target data and learned knowledge to avoid catastrophic forgetting.
-DSM selectively updates each AC based on its output reliability, thereby alleviating reliability bias.
-The SPS enhances the model’s robustness to label noise by assigning varying probabilities to multiple labels, which mitigates potential error accumulation.
-contribution of thiss work can be summarised as follows:
-1. We c
-1). We propose a method, MDAA, for a more challenging TTA task named MM-CTTA, and explain why typical TTA methods are not well-suited for this task, as illustrated by a example in Fig.1.
-2).
-We innovatively apply AC to TTA to keep the model from catastrophic forgetting during adaption.
-We propose DSM and SPS to further dynamically integrate the features of different modal thus mitigating reliability bias the error accumulation.
-3).
-We design two MM-CTTA tasks to meet the needs of real-world environments.
-Extensive experiments demonstrate that MDAA achieves SOTA performance, surpassing previous methods by up to 6.22% and 6.84% in the two tasks respectively.
-2 Related Works
-2.1 Test-Time Adaptation
-Test-Time Adaptation (TTA) focuses on enabling a pre-trained model to adapt to a new target domain without requiring access to the source domain data used to initially train the model.
-A major challenge in TTA is error accumulation, which affects methods that rely on pseudo-labeling; incorrect predictions can mislead the adaptation process, potentially causing the model to collapse (Chen et al., 2019).
-One of the early solutions, TENT, addresses this by only updating the model’s batch normalization (BN) layers through entropy minimization (Wang et al., 2021).
+Abstract.
+Test-Time Adaptation (TTA) enables pre-trained models to bridge the gap between source and target datasets using unlabeled test data, addressing domain shifts caused by corruptions like weather changes, noise, or sensor malfunctions in test time.
+Multi-Modal Continual Test-Time Adaptation (MM-CTTA), as an extension of standard TTA, further allows models to handle multi-modal inputs and adapt to continuously evolving target domains.
+However, MM-CTTA faces critical challenges such as catastrophic forgetting and reliability bias, which are rarely addressed effectively under multi-modal corruption scenarios.
+In this paper, we propose a novel approach, Multi-modality Dynamic Analytic Adapter (MDAA), to tackle MM-CTTA tasks.
+MDAA introduces analytic learning—a closed-form training technique—through Analytic Classifiers (ACs) to mitigate catastrophic forgetting.
+Furthermore, we design the Dynamic Late Fusion Mechanism (DLFM) to dynamically select and integrate reliable information from different modalities.
+Extensive experiments show that MDAA achieves state-of-the-art performance across the proposed tasks.
+1. Introduction
+Test-Time Adaptation (TTA) aims to enable pre-trained models to bridge the gap between the source domain and the target domain (Wang et al., 2021; Liang et al., 2024).
+Unlike Unsupervised Domain Adaptation (UDA) (Zhang et al., 2015), TTA performs adaptation without any source data (i.e., pre-trained dataset), which not only saves computational resources by avoiding retraining but also preserves the privacy of the source data.
+One key application of TTA is to address the problem of domain shift from source data to corrupted test data, where the corruption is often caused by weather changes, ambient noise or sensor malfunctions.
+In these scenarios, Multi-Modal Test-Time Adaptation (MM-TTA) offers a distinct advantage over mono-modal TTA, as domain shifts often affect only a subset of modalities, while others remain relatively stable and informative (Radford et al., 2021; Yang et al., 2024).
+This allows the model to leverage uncorrupted modalities to guide adaptation, thereby enhancing robustness in complex real-world environments.
+To better reflect real-world settings where distribution shifts are continuous and multi-faceted, Multi-Modal Continual Test-Time Adaptation (MM-CTTA) has been proposed (Cao et al., 2023).
+MM-CTTA addresses more complicated domain shifts, where the target distribution evolves both over time and across modalities.
+This setting is common in real-world scenarios, particularly for general-purpose AI agents (e.g., autonomous driving, robotic systems) that rely on multiple sensors for perception, as illustrated in Figure 1.A.
+However, achieving reliable MM-CTTA remains challenging, mainly due to two key issues.
+The first is reliability bias, which arises from intra-modality domain shifts and exacerbates inconsistencies within downstream fusion networks (Yang et al., 2024).
+The second is catastrophic forgetting, where knowledge acquired from the source domain deteriorates during continual adaptation, ultimately reducing the model’s generalization capability (McCloskey and Cohen, 1989).
+MM-CTTA models must continuously adapt to multi-modal domain shifts while mitigating modality collapse and ensuring reliability in real-world scenarios.
+While recent researches on TTA have achieving promising results of in catastrophic forgetting (Wang et al., 2022; Gao et al., 2023; Niu et al., 2022) and reliability bias (Yang et al., 2024; Lei and Pernkopf, 2024) respectively, directly combining CTTA methods with MM-TTA works often falls short.
+This is because dynamically changing corruption patterns magnify reliability bias and create additional complexities, as corruption can unpredictably switch between modalities during continual adaptation.
+Methods explicitly designed for MM-CTTA remain scarce.
+The closest existing approach, CoMAC (Cao et al., 2023), focuses primarily on segmentation tasks in dynamic environments rather than addressing modality corruption.
+In this paper, we propose a new approach named Multi-modality Dynamic Analytic Adaptor (MDAA) to address the challenges in MM-CTTA.
+MDAA mainly comprises two primary components: (i) the Analytic Classifiers (ACs), and (ii) the Dynamic Late Fusion Mechanism (DLFM).
+The ACs update the model by addressing a recursive ridge regression problem, optimizing on both new target data and learned knowledge to avoid catastrophic forgetting, such approach have already been used in continual learning (CL) field (Zhuang et al., 2022, 2023).
+Moreover, each feature type in the multi-modal model is configured with an individual AC in our framework.
+As a result, MDAA can derive predictions not only from the downstream fusion network but also from the feature encoders, marking a significant departure from conventional multi-modal paradigms.
+To effectively utilize outputs from different ACs, we proposed a tailored late fusion strategy, DLFM, which compensates for and integrates predictions across modalities.
+DLFM selects reliable predictions from modality-specific ACs and propagates them to guide the updates of ACs associated with corrupted modalities, thereby alleviating reliability bias.
+To rigorously assess the robustness and adaptability of MDAA, we design two challenging evaluation settings: (1) adaptation over extremely long phases and (2) scenarios with alternating modality corruption.
+These settings provide a rigorous test of MDAA’s robustness and adaptability under real-world conditions.
+The key contributions of this work are listed as follows:
+1).
+We introduce ACs to TTA into the TTA framework to prevent the model from catastrophic forgetting.
+We propose DLFM to adaptively fuse features from multiple modalities, thereby mitigating reliability bias.
+2). We design two MM-CTTA benchmarks featuring modality corruption switching and prolonged adaptation, better capturing real-world dynamics, pushing the boundaries of MM-CTTA.
+3). MDAA achieves state-of-the-art performance, outperforming prior methods by up to 6.22% and 6.84% in two benchmarks, while being more efficient in runtime and memory usage.
+2. Related Works
+2.1. Test-Time Adaptation
+TTA focuses on enabling a pre-trained model to adapt to a new target domain without requiring access to the source domain data.
+However, simply relying on incorrect predictions during adaptation can mislead the update process, potentially leading to model collapse (Chen et al., 2019).
+TENT, as one of the first methods solve such issue, addresses this by only updating the model’s batch normalization (BN) layers through entropy minimization (Wang et al., 2021).
 Subsequent research (Niu et al., 2023; Gong et al., 2022a; Zhou et al., 2023) has further mitigated this issue by filtering out low-confidence predictions using carefully designed thresholds and updating the layer normalization (LN) layers for more robust performance.
-2.2 Continual Test-Time Adaptation
-Continual Test-Time Adaptation (CTTA) extends TTA to scenarios where the target domain changes continuously in an online manner without access to source data.
-The additional challenge CTTA faces is known as catastrophic forgetting, which occurs when a model adapts to the target domain, leading to the loss of knowledge acquired from the source domain and dimension of generalization ability.
-To solve this problem, some studies turn to Continual Learning (CL) and achieve great success (Niu et al., 2022; Cao et al., 2023).
-We follow this trend and introduce a novel CL approach called Analytic Continual Learning (ACL) in MDAA.
-ACL provides a global optimal solution through matrix inverse operations (Guo & Lyu, 2004).
-To address the out-of-memory issue caused by large inverse matrices, Zhuang et al.
-(2021a) demonstrated that iterative computation using block-wise data achieves results equivalent to joint computation, making analytic learning highly effective in continual learning.
-By treating data from different time periods as blocks, ACL allows for recursive computation, with the final result being as accurate as if all data were processed simultaneously.
-Thanks to its non-forgetting properties, ACL has shown strong performance across various CL tasks in recent years (Zhuang et al., 2022; 2023).
-Inspired by ACL’s success in CL tasks, we apply ACL to TTA for the first time in this work, implementing several enhancements to address the issue of catastrophic forgetting.
-2.3 Multi-modality Test-Time Adaptation
-Multi-modality Test-Time Adaptation (MM-TTA) seeks to enhance model reliability by incorporating multi-modal data into the TTA task.
-However, imbalances in inter-modal reliability can result in significant performance degradation, known as reliability bias (Wang et al., 2020).
-Most existing MM-TTA models address this issue by independently updating BN or LN layers of each feature encoder, followed by a weighted fusion mechanisms (Shin et al., 2022; Cao et al., 2023; Xiong et al., 2024).
+2.2. Multi-modality Test-Time Adaptation
+MM-TTA (Zhao et al., 2025; Guo and Jin, 2025; Yang et al., 2024; Wang et al., 2020) seeks to enhance model reliability by incorporating multi-modal data into the TTA task.
+However, imbalances in inter-modal reliability can result in significant performance degradation, known as reliability bias (Wang et al., 2020), shown in Figure 1.B.
+Existing MM-TTA models address this issue by independently updating BN or LN layers of each feature encoder, followed by a weighted fusion mechanisms (Shin et al., 2022; Cao et al., 2023; Xiong et al., 2024).
 Although this allows more reliable modalities to carry more weight during fusion, the approach remains relatively shallow in terms of information integration.
@@
 READ achieves reliable adaptation by modulating only the fusion layer within the attention module of the ViT block.
-Follow-up work (Lei & Pernkopf, 2024) aimed to improve performance by updating both the feature encoders and the fusion layer.
+Follow-up work (Lei and Pernkopf, 2024) aimed to improve performance by updating both the feature encoders and the fusion layer.
 However, this approach requires prior knowledge about which modalities are corrupted, making it less capable for real-world scenario.
-In this paper, our MDAA approach introduces classifiers for each feature encoder and the fusion network.
-By adjusting the parameters of each classifier individually, MDAA maximizes the use of information from each modality, thus mitigating the effects of reliability bias.
-Crucially, these classifier updates for upstream and downstream blocks do not cause conflicts, as the entire pre-trained model remains frozen throughout the process.
-This property enables MDAA to adapt to changing modality corruption without extra prior knowledge, distinguishing it from existing MM-TTA approaches.
-3 Method
-In this section, we first define the challenging MM-CTTA setting and introduce notations for key concepts in Sec.
-3.1.
-We then introduce the proposed Multi-modality Dynamic Analytic Adapter (MDAA) approach, which includes Analytic Classifiers (ACs) integrated with pre-trained multi-modal encoders (Sec.
-3.2), the Dynamic Selection Mechanism (DSM) (Sec.
-3.3), and the Soft Pseudo-label Strategy (SPS) (Sec.
-3.4).
-An overview of MDAA is illustrated in Fig.
-2, and the pseudo-code for MDAA is provided in Appendix B.
-3.1 Problem Definition and Notations
-In this paper, we focus on an audio-video classification task as an example, using two modalities for illustration without loss of generality.
-In MM-CTTA, the pre-trained model is trained on a labeled source dataset in source domain S, where and represent the audio and video training data matrices respectively.
+2.3. Analytic Continual Learning
+When extending TTA to a continual setting, a significant challenge that emerges is catastrophic forgetting—a phenomenon where adaptation to the target domain leads to degradation of knowledge learned from the source domain and compromises generalization (as illustrated in Figure 1.C).
+Recent works have explored the integration of Continual Learning (CL) into TTA, demonstrating promising results (Niu et al., 2022; Cao et al., 2023).
+Following this line of research, we incorporate a novel CL paradigm, Analytic Continual Learning (ACL), into our proposed MDAA framework.
+ACL provides a global optimal solution through matrix inverse operations (Guo and Lyu, 2004).
+To address the out-of-memory issue caused by large inverse matrices, (Zhuang et al., 2021) demonstrated that iterative computation using block-wise data achieves results equivalent to joint computation, making analytic learning highly effective in continual learning.
+By treating data from different time periods as blocks, ACL allows for recursive computation, with the final result being as accurate as if all data were processed simultaneously.
+Thanks to its non-forgetting properties, ACL has shown strong performance across various CL tasks in recent years (Zhuang et al., 2022, 2023).
+3. Multi-modality Dynamic Analytic Adapter
+In this section, we first formalize the MM-CTTA problem and introduce the key notations in Sec.3.1.
+Then present the proposed MDAA framework, which incorporates ACs into pre-trained multi-modal encoders (Sec.3.2).
+Subsequently, we detail the AC adaptation in Sec.3.3 and introduce DLFM for updating process in Sec.3.4.
+An overview of MDAA is provided in Figure 2.
+3.1. Problem Definition and Notations
+To illustrate our method, we consider an audio-video classification task, which generalizes to other multi-modal scenarios without loss of generality.
+In MM-CTTA, the weight in pre-trained models is trained on a labeled source dataset in source domain Sr, where and represent the audio and video training data respectively.
 represents the corresponding one-hot label set.
-During the adaptation phase, for each timestamp in the target domain T, the model must perform inference and update its parameters based on the unlabeled test dataset .
-The suffix indicates the current target domain, as it shifts continually.
-Note that only the test dataset at timestamp is available for updating the parameters .
-3.2 Source Model and Analytic Classifier Configuration
-In the context of multi-modality classification task, we propose to integrate the ACs as classifiers into a typical extraction-fusion approach. The standard structure can be represented as follows:
-where and represent upstream feature encoders for audio and video modalities, respectively, indicates the fusion operation implemented as concatenation in this paper, denotes the downstream fusion network and indicates the classifier.
-Specifically, we leverage multiple ACs as classifiers for extracted features of each modality and the fused features, with each classifier making an independent prediction.
-Each AC is of the same structure as a two-layer fully connected network, denoted by .
-Features are first non-linearly projected into a higher dimensional space by the feature expansion layer , enhancing their expressiveness (Zhuang et al., 2022; 2021b).
-The projected features are then passed through the linear layer for classification.
-The feature expansion layer remains frozen, while the linear layer requires updates during the adaptation process.
-For simplicity, we combine the feature encoder and the feature expansion layer into a single function, denoted as , and refer to the projected features as .
-Consequently, MDAA yields three classifiers as follows:
-Since all ACs can be represented by the same equation, we will not distinguish between them except for special needs in the following discussion.
-Following previous ACL works Zhuang et al.
-(2022), the classifier are updated by solving the ridge regression to optimize on the source data .
-To further solve the class imbalance problem within the source data, inspired by Fang et al.
-(2024), we formulate the optimization problem as follows:
+During the adaptation phase, for each timestamp in the target domain Tg, the model can only perform inference and update itself based on the unlabeled test dataset .
+The target domain varies with and the whole adaptation process can be represented as .
+3.2. Source Model and Analytic Classifier Configuration
+The multi-modal backbone follows a typical extraction–fusion architecture, where training data from different modalities are first processed by modality-specific feature encoders and subsequently fused through a fusion encoder.
+During this process, features from different encoders are projected into a unified higher-dimensional space to enhance representational capacity, and are then passed through a linear classifier to produce prediction labels, as illustrated in Figure 2.A.
+These linear classifiers are referred to as Analytic Classifiers (ACs).
+During the adaptation phase, both the feature encoders and the projection layers are kept frozen, while only the ACs are updated.
+Since features from different modalities are processed by ACs in a unified manner, we omit modality-specific distinctions for simplicity.
+The weights for the ACs are optimized via ridge regression, initialized using a least squares objective, a strategy commonly employed in classification tasks (Zhuang et al., 2022; Fang et al., 2024; Zhuang et al., 2023; Yang and Hu, 2008).
+To address the issue of class imbalance, we reformulate the optimization objective as follows:
 where indicates the Frobenius norm, is the regularization parameter and is the samples size of .
 , and represents the weight, expanded feature vector and one-hot label of sample k in , while the weight is further defined as
-where is the number of classes in and is the number of samples in from category to which sample belongs.
-Following the ridge regression solution, the solution to optimization problem is given in Theorem 1.
-Theorem 1. The optimal solution to Formula 4 is given as
+where is the number of classes in and is the number of samples in from category to which sample belongs. Following the ridge regression solution, the solution to optimization problem is given as
 where and .
-The proof of Theorem 1 is provided in Appendix A.
 In addition to the classifier weights , a memory bank needs to be constructed during the training phase.
 Unlike other methods (Cao et al., 2023; Zhang et al., 2023; Xiong et al., 2024), our memory bank contains only two types of matrices, which can be represented as , where
-Both and are used to extract and preserve the learned knowledge from the source dataset, which cannot be accessed during adaptation. Therefore weight can be further rewritten as
-3.3 Dynamic Selection Mechanism
-When adapting to sequentially incoming multi-modal data in MM-CTTA, some modality may be unreliable due to corruption.
-Such corrupted data can mislead the model to learn incorrect information.
-To address the reliability bias, we propose Dynamic Selection Mechanism (DSM) to determine whether each AC should be updated in a dynamic way.
-M first identifies the most reliable classifier among the three (i.e., a, v, m) as the leader, while the other classifiers are treated as follower.
-Specifically, the leader is determined by comparing the maximum probability from each classifier’s distribution.
-For each sample, the model prediction corresponds to the leader’s prediction, and the leader will not be updated to maintain class balance.
-Whether to update each follower in the model depends on the comparison between its maximum probability distribution and the leader’s maximum probability distribution .
-We consider four possible scenarios, as illustrated in Fig.3:
+Both and are used to extract and preserve the learned knowledge from the source dataset, which cannot be accessed during adaptation. Therefore can be further rewritten as
+3.3. Model Inference and Analytic Classifier Adaptation
+During test time, MDAA determines the final prediction for each sample by selecting the output distribution with the highest maximum a posteriori (MAP) score across all ACs, as higher predictive probabilities generally reflect greater confidence (Boudiaf et al., 2022).
+Subsequently, DLFM is employed to further filter the samples used for model updates, which will be elaborated upon in the following section.
+Since ACL performs global optimization, it considers not only the currently selected input sample at timestamp , but also all previously observed data, including and .
+The corresponding optimization problem for the weight is formulated as:
+where denotes for soft pseudo-labels of filtered samples.
+We adopt soft labels for AC updates as they retain a richer representation of uncertainty compared to hard labels (Müller et al., 2019; Hinton et al., 2015).
+For each filtered sample, we sort and choose its top- class positions and assign weights .
+In this paper, we define the weights as , and construct the soft pseudo-label accordingly as:
+Such reconsrtuction is proven to be useful, as shown in ablation study in subsection 4.3.
+Due to the definition of TTA, all datasets (i.e., and ) prior to timestamp are not accessible when solving the optimization problem.
+However, with the aid of the memory bank , the solution can still be computed, as stated follow.
+Proof of Theorem 1 is provided in Appendix D.
+It can be seen that, only features and labels that are accessible in phase are involved in the adaptation, which fits the definition of TTA.
+Solution in Eq.
+8 is the global optimal solution for all the samples that have appeared, therefore it is not affected by batch size and timestamp, thus solving the catastrophic forgetting at the root.
+3.4. Dynamic Late Fusion Mechanism
+Although AC-based adaptation effectively mitigates the forgetting issue, it is worth noting that employing the least squares loss as the objective function may be less robust to outliers compared to the cross-entropy loss.
+This vulnerability increases the risk of model collapse when adapting with corrupted modality inputs.
+This raises a critical question: how can we reliably select trustworthy samples for updating each AC?
+To tackle this challenge, we propose a novel fusion strategy termed Dynamic Late Fusion Mechanism (DLFM).
+For each input sample, DLFM first identifies the AC yielding the highest MAP score, designating it as the Leader.
+During adaptation, each AC is updated independently based on its MAP score relative to the Leader.
+All possible relationships and update rules can be enumerated as four scenarios (see Figure 3).
 (i).
-Close Distributions: and are quite close and refer to the same label.
-In this case, the follower is not updated, as this would only reinforce what it has already learned, potentially leading to an imbalanced class distribution.
+Close Distributions: When the MAP scores of the AC and the Leader are nearly identical and correspond to the same label, the AC remains unchanged to avoid reinforcing existing knowledge and causing class imbalance.
 (ii).
-Different Labels with Close Probabilities: and are quite close while they refer to different labels.
-The follower is also not updated in this scenario because there is no certainty that the leader’s result is correct, and updating may introduce errors.
-(iii). Evenly Distributed Probabilities: Both the leader and follower have evenly distributed probabilities with no significant difference between labels. Again, no update occurs.
-(iv). Significant Difference: The leader has a higher probability label, while the follower has a more even probability distribution. In this scenario, the follower should be updated.
-In general, cases (i),(ii) and (iii) all belong to a small gap between and , while case (iv) belongs to a larger gap. Therefore, given a pre-defined threshold , DSM can be noted as:
-3.4 Soft Pseudo-label Strategy
-Using soft labels in self-supervised learning is quite popular as it preserves a broader range of possible outcomes compared to hard (one-hot) label learning (Müller et al., 2019; Hinton et al., 2015).
-Inspired by such trend, we use soft pseudo-labels to update the ACs during MM-CTTA, as illustrated in Fig.2(C).
-For each test sample , we choose the top classes of leader’s distribution set and assign them with weights , () respectively.
-The reconstructed label for sample through SPS can be represented as
-Since ACL considers global optimization, which means it accounts not only for the input data at timestamp , but also all previous data processed by the model, including and .
-With the reconstructed label throughout , the optimization problem for weight matrix can be represented as:
-It is important to note that while we treat source data and target data separately, we adjust the category balance for the source dataset by assigning weights to each class.
+Different Labels with Close Distributions: If the MAP scores are similar but correspond to different labels, no update is made to avoid introducing potential errors due to uncertainty in the Leader’s prediction.
+(iii).
+Evenly Distributed Probabilities: If both the AC and the Leader show evenly distributed probabilities with no clear preference, the AC is not updated to prevent adjustments based on uncertain predictions.
+(iv). Significant Difference: If the Leader’s MAP is significantly higher than the AC’s, the AC is updated to reflect the Leader’s confident prediction.
+To summarize, cases (i),(ii) and (iii) all belong to a small gap between MAP scores of AC and Leader, while case (iv) belongs to a larger gap.
+Therefore, given a pre-defined threshold , Rules of DLFM can be noted as:
+DLFM leverages the complementary and consistent information across modalities to enhance adaptation.
+For complementarity, it selects the most reliable prediction—typically from an uncorrupted modality—as a pseudo label to update other classifiers, aligning corrupted inputs with reliable targets to handle domain shifts.
+For consistency, when all classifiers predict the same label, indicating high confidence and low likelihood of corruption, no update is performed to avoid unnecessary adaptation and prevent class imbalance.
+Since our recursive formulation already preserves prior knowledge, additional updates are unnecessary.
+By updating each classifier individually, DLFM maximizes modality-specific information while reducing reliability bias.
+It is also important to note that in Eq.
+3.3 we treat source data and target data separately, we adjust the category balance for the source dataset by assigning weights to each class.
 In contrast, it is challenging to apply this approach to the target dataset, as it is impossible to know the exact number of samples in each category within the target domain.
-However, with the help of the DSM, category balance can still be maintained, as unimportant samples have been filtered out.
+However, with the help of the DLFM, category balance can still be maintained, as unimportant samples have been filtered out.
 Given that the weights added to the sample data are expected to average out to 1, the impact of each sample from both the source and target domains on the model can be considered equivalent.
-Due to the definition of TTA, all datasets (i.e., and ) prior to timestamp are not accessible when solving the optimization problem.
-However, with the aid of the memory bank , the solution can still be computed, as stated in Theorem 2.
-Theorem 2. The optimal solution to Formula 12 is given as
-where the memory bank is updated in a recursive way with timestamp , as
-The proof of Theorem 2 is also provided in Appendix A.
-It can be seen that, the size of the memory bank is constant and depends on the dimension of the feature expansion layer and class number .
-The equations above demonstrate that the size of remains unchanged during adaptation, regardless of the timestamp .
-4 Experiments
-In this section, we evaluate the proposed MDAA approach under the challenging MM-CTTA setting.
-The MM-CTTA setting is detailed in Sec.
-4.1.
-Sec.
-4.2 compares MDAA with SOTA methods through extensive experiments to demonstrate its superior performance.
-Additionally, the ablation studies on each component of MDAA are presented in Sec.
-4.3, which illustrate the effectiveness of MDAA in addressing different challenges within the MM-CTTA setting.
-The implementation details are provided in Appendix C.
-4.1 Benchmarks and Settings
-In this section, we introduce the datasets and task settings used for MM-CTTA.
-The MM-CTTA setting requires the model to initially train on uncorrupted source datasets.
-Subsequently, the model performs TTA on each corrupted target domain in sequence.
-We utilize two datasets for this setting: Kinetics50 (Kay et al., 2017) and VGGSound (Chen et al., 2020).
-While the original uncorrupted datasets serve as source, the corrupted target datasets, Kinetics50-C and VGGSound-C, are constructed following Yang et al.
-(2024), which introduces 15 types of video corruptions and 6 audio corruptions at severity level 5.
-To evaluate the model performance, we designed two classification tasks specifically for MM-CTTA following previous research on corruption-related TTA (Wang et al., 2022; 2021; Yang et al., 2024; Lei & Pernkopf, 2024).
-The first task, named progressive single-modality corruption, sequentially introduces different types of corruption to one modality while keeping the other modality uncorrupted.
-Focusing on evaluating the model’s resistance to catastrophic forgetting, this task is set in an online manner, where the model processes only one sample at a time.
-The second task, called interleaved modality corruption, continually alternates corruption between the two modalities.
-While most methods perform poorly in the online setting due to severe catastrophic forgetting, this task uses a batch size of 64 during test time to emphasize assessing the model’s ability to adapt to dynamic reliability biases.
-4.2 Performance Comparison
-To provide a comprehensive comparison, we reproduce different types of TTA methods under the MM-CTTA setting.
-Typical TTA methods of TENT (Wang et al., 2021) and SAR (Niu et al., 2023); CTTA methods of CoTTA (Wang et al., 2022) and EATA (Niu et al., 2022); and MM-TTA methods of MMTTA (Shin et al., 2022) and READ (Xiong et al., 2024).
-To ensure a fair comparison, all methods are based on the pre-trained ViT-baesed CAV-MAE (Gong et al., 2022b) as the multi-modal encoders.
-When reproduce methods that update the BN layers, we instead update the LN layers to suit the ViT structure.
-Additionally, we evaluate the Source model as a strong baseline, which is only trained on the source dataset and remains frozen during test time.
-Performances of different methods on progressive single-modality corruption are listed in Table 1, 2 and 3, where Table 1 is based on the audio corruption, Table 2 and 3are performance on the video corruption in Kinetics50-C and VGGSound-C.
-The proposed MDAA achieves SOTA performance on most target domains.
-It is noteworthy that nearly all comparison models collapse and perform significantly worse than non-updating models on VGGSound-C.
-In contrast, EATA achieves better results by restricting model parameter updates, which facilitates successful adaptation.
-Specifically, MDAA outperforms previous methods by 3.00%-3.57% and 3.03%-6.22% on average for audio and video tasks in Kinetics50-C, and by 0.94%-1.10% and 0.13%-0.18% on average for audio and video tasks in VGGSound-C.
-Furthermore, MDAA consistently maintains its lead over other methods in the later stages of the tasks.
-These results demonstrate that MDAA is more robust against catastrophic forgetting in MM-CTTA.
 | Method | Type | Kinetics50-C | VGGSound-C | ||||||||||||
@@
 | SAR | TTA | 72.18 | 70.36 | 48.30 | 37.67 | 36.21 | 39.09 | 50.64 | 16.09 | 4.50 | 4.33 | 3.60 | 12.00 | 5.51 | 7.67 |
-| CoTTA | CTTA | 19.67 | 4.10 | 2.11 | 2.03 | 2.03 | 2.03 | 5.33 | 5.85 | 1.35 | 0.52 | 0.53 | 0.57 | 0.38 | 1.53 |
 | EATA | CTTA | 73.91 | 65.29 | 68.24 | 70.51 | 68.28 | 70.48 | 69.45 | 40.39 | 31.99 | 31.91 | 32.38 | 39.24 | 33.95 | 34.98 |
 | MMTTA* | MM-TTA | 17.03 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 4.50 | 0.41 | 0.33 | 0.33 | 0.33 | 0.33 | 0.33 | 0.34 |
-| READ | MM-TTA | 68.33 | 59.75 | 57.38 | 54.14 | 53.49 | 52.72 | 57.63 | 18.53 | 7.99 | 7.44 | 5.71 | 8.19 | 4.73 | 8.77 |
+| READ | MM-TTA | 65.65 | 49.34 | 46.45 | 44.94 | 44.45 | 43.72 | 49.09 | 18.73 | 8.52 | 8.01 | 5.90 | 8.30 | 4.82 | 9.05 |
+| READ+SRM | MM-CTTA | 52.23 | 38.30 | 40.68 | 46.61 | 47.02 | 48.60 | 45.57 | 18.36 | 7.88 | 8.34 | 8.01 | 16.26 | 10.21 | 11.51 |
 | MDAA | MM-CTTA | 72.87 | 71.45 | 72.91 | 72.26 | 73.20 | 73.80 | 72.75 | 38.80 | 34.91 | 34.63 | 34.59 | 37.70 | 35.85 | 36.08 |
@@
 | SAR | TTA | 41.81 | 27.94 | 24.41 | 40.47 | 42.90 | 70.36 | 41.32 | 14.91 | 4.56 | 4.61 | 3.72 | 12.44 | 5.94 | 7.70 |
-| CoTTA | CTTA | 1.99 | 1.99 | 2.92 | 5.07 | 11.64 | 32.56 | 9.36 | 0.30 | 0.30 | 0.30 | 0.30 | 0.30 | 0.36 | 0.31 |
 | EATA | CTTA | 73.91 | 65.32 | 68.18 | 70.49 | 68.26 | 70.47 | 69.44 | 40.22 | 33.69 | 31.61 | 32.64 | 39.67 | 32.81 | 35.11 |
 | MMTTA* | MM-TTA | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 39.21 | 8.19 | 0.33 | 0.30 | 0.31 | 0.30 | 0.61 | 1.49 | 0.56 |
-| READ | MM-TTA | 56.50 | 56.09 | 56.30 | 57.25 | 62.99 | 65.14 | 59.05 | 9.20 | 5.82 | 7.48 | 7.89 | 11.67 | 12.11 | 9.03 |
+| READ | MM-TTA | 48.95 | 48.70 | 48.72 | 51.99 | 62.69 | 65.34 | 54.40 | 8.99 | 6.00 | 7.71 | 7.71 | 11.56 | 11.89 | 8.98 |
+| READ+SRM | MM-CTTA | 48.59 | 47.39 | 46.48 | 41.11 | 38.01 | 51.90 | 45.58 | 18.50 | 7.93 | 8.36 | 7.96 | 16.28 | 10.11 | 11.53 |
 | MDAA | MM-CTTA | 74.86 | 72.63 | 72.87 | 72.26 | 73.4 | 72.02 | 73.01 | 38.95 | 35.46 | 34.66 | 34.70 | 37.31 | 35.20 | 36.05 |
@@
 | SAR | TTA | 38.36 | 35.97 | 34.51 | 44.40 | 48.86 | 50.77 | 47.53 | 43.59 | 35.81 | 42.54 | 52.11 | 35.44 | 50.20 | 40.15 | 50.73 | 43.40 |
-| CoTTA | CTTA | 33.43 | 27.51 | 25.20 | 21.19 | 18.19 | 16.41 | 14.91 | 13.29 | 11.18 | 9.60 | 8.43 | 6.89 | 6.36 | 5.39 | 4.09 | 14.80 |
 | EATA | CTTA | 48.80 | 49.82 | 49.03 | 67.66 | 61.98 | 70.84 | 66.16 | 61.64 | 61.54 | 45.40 | 75.99 | 51.95 | 65.88 | 68.71 | 66.08 | 60.77 |
 | MMTTA* | MM-TTA | 14.31 | 2.64 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.89 |
-| READ | MM-TTA | 11.92 | 2.04 | 2.03 | 2.97 | 2.41 | 2.46 | 2.41 | 2.30 | 2.04 | 2.04 | 2.04 | 2.04 | 2.04 | 2.04 | 2.03 | 2.86 |
+| READ | MM-TTA | 12.76 | 2.03 | 2.02 | 3.27 | 2.60 | 2.64 | 2.64 | 2.50 | 2.39 | 2.42 | 2.44 | 2.40 | 2.42 | 2.40 | 2.41 | 3.16 |
+| READ+SRM | MM-CTTA | 41.90 | 42.29 | 42.27 | 54.78 | 54.37 | 58.21 | 56.70 | 53.70 | 52.68 | 46.12 | 61.97 | 45.16 | 58.63 | 55.32 | 54.54 | 52.68 |
 | MDAA | MM-CTTA | 54.89 | 55.25 | 55.32 | 63.89 | 62.49 | 67.26 | 65.86 | 64.32 | 65.31 | 61.86 | 73.20 | 61.60 | 67.83 | 69.22 | 68.69 | 63.80 |
@@
 | SAR | TTA | 34.75 | 35.08 | 35.89 | 42.70 | 45.99 | 49.43 | 50.12 | 44.08 | 42.42 | 40.02 | 57.54 | 35.56 | 48.86 | 57.22 | 66.26 | 45.73 |
-| CoTTA | CTTA | 3.89 | 4.01 | 4.50 | 5.39 | 5.47 | 5.75 | 7.82 | 4.98 | 5.79 | 9.85 | 6.89 | 12.72 | 14.26 | 22.33 | 51.34 | 11.00 |
 | EATA | CTTA | 48.81 | 49.79 | 49.02 | 67.71 | 61.96 | 70.88 | 66.17 | 61.56 | 61.51 | 45.38 | 75.96 | 51.90 | 65.90 | 68.76 | 66.09 | 60.76 |
 | MMTTA* | MM-TTA | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 1.99 | 23.88 | 3.45 |
-| READ | MM-TTA | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.19 | 2.93 | 22.95 | 3.50 |
+| READ | MM-TTA | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.03 | 2.09 | 2.03 | 2.22 | 2.04 | 2.43 | 3.65 | 24.34 | 3.67 |
+| READ+SRM | MM-CTTA | 41.91 | 42.42 | 42.46 | 54.97 | 54.32 | 58.30 | 56.83 | 53.66 | 52.93 | 46.03 | 62.05 | 45.00 | 58.69 | 55.20 | 54.60 | 53.19 |
 | MDAA | MM-CTTA | 67.32 | 67.48 | 67.76 | 68.98 | 67.60 | 69.59 | 68.49 | 66.46 | 66.18 | 63.14 | 72.87 | 59.33 | 66.59 | 67.64 | 65.25 | 66.98 |
-The comparison results of interleaved modality corruption tasks are shown in Table 4 and 5.
-In this task, EATA, which is more good at memorisation, is not dominant in the task of highlighting reliability bias.
-READ, which is specifically designed to address intra-modal reliability bias, demonstrates strong performance in this area.
-However, its effectiveness is limited to video corruption in MM-CTTA, as performance drops significantly during audio corruption.
-In contrast, MDAA is well-adapted to the corruption of different modalities, outperforming READ by 2.39%-6.84% and EATA by 0.60%-7.28% on average across Kinetics50-C and VGGSound-C.
 | Method | Type | Gauss. | Shot. | Impul. | Defoc. | Glass. | Motion. | Zoom. | Snow | Frost | Fog | Bright. | Cont. | Elastic. | Pixel. | Jpeg | Avg. |
@@
 | SAR | TTA | 43.12 | 38.99 | 37.77 | 42.43 | 43.84 | 43.61 | 43.79 | 42.13 | 41.26 | 42.83 | 43.84 | 39.34 | 42.75 | 43.49 | 40.38 | 41.97 |
-| CoTTA | CTTA | 31.20 | 6.93 | 0.51 | 0.38 | 1.26 | 0.95 | 0.45 | 0.33 | 0.33 | 0.33 | 0.27 | 0.32 | 0.34 | 0.36 | 0.40 | 2.96 |
 | EATA | CTTA | 53.57 | 53.70 | 53.57 | 57.00 | 57.29 | 58.46 | 57.77 | 56.24 | 56.57 | 55.45 | 59.06 | 54.13 | 58.24 | 57.22 | 57.38 | 56.38 |
 | MMTTA* | MM-TTA | 0.46 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.35 |
-| READ | MM-TTA | 33.02 | 0.33 | 0.33 | 0.41 | 0.33 | 0.33 | 0.33 | 0.33 | 0.33 | 0.33 | 0.33 | 0.33 | 0.33 | 0.33 | 0.33 | 2.51 |
+| READ | MM-TTA | 33.12 | 10.41 | 4.05 | 0.77 | 0.37 | 0.35 | 0.32 | 0.29 | 0.29 | 0.29 | 0.29 | 0.29 | 0.29 | 0.29 | 0.29 | 3.45 |
+| READ+SRM | MM-CTTA | 41.33 | 41.37 | 41.32 | 45.50 | 45.68 | 47.18 | 46.61 | 44.66 | 44.88 | 42.97 | 47.07 | 40.01 | 46.18 | 45.41 | 45.24 | 44.36 |
 | MDAA | MM-CTTA | 55.13 | 55.29 | 55.30 | 56.91 | 57.20 | 57.78 | 57.32 | 56.52 | 56.25 | 56.14 | 58.11 | 55.32 | 57.06 | 56.27 | 57.39 | 56.53 |
@@
 | SAR | TTA | 40.31 | 39.24 | 38.33 | 41.53 | 41.36 | 44.43 | 43.59 | 42.46 | 41.11 | 41.52 | 42.24 | 38.97 | 43.49 | 40.97 | 46.22 | 41.72 |
-| CoTTA | CTTA | 0.33 | 0.33 | 0.35 | 0.33 | 0.33 | 0.33 | 0.32 | 0.46 | 0.36 | 0.35 | 0.56 | 2.24 | 8.36 | 19.22 | 38.95 | 4.85 |
 | EATA | CTTA | 53.63 | 53.60 | 53.61 | 57.06 | 57.27 | 58.35 | 57.83 | 56.22 | 56.74 | 55.73 | 59.16 | 54.09 | 58.19 | 57.27 | 57.35 | 56.41 |
 | MMTTA* | MM-TTA | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 0.34 | 8.32 | 0.87 |
-| READ | MM-TTA | 0.32 | 0.32 | 0.32 | 0.32 | 0.32 | 0.32 | 0.32 | 0.32 | 0.89 | 16.32 | 25.48 | 26.95 | 32.93 | 40.25 | 50.31 | 13.05 |
+| READ | MM-TTA | 13.26 | 13.39 | 13.38 | 13.66 | 13.72 | 13.91 | 13.92 | 14.25 | 14.84 | 18.15 | 20.68 | 22.85 | 29.46 | 40.99 | 50.38 | 20.46 |
+| READ+SRM | MM-CTTA | 41.46 | 41.25 | 41.41 | 45.77 | 45.95 | 47.17 | 46.85 | 44.73 | 45.03 | 43.37 | 46.90 | 39.93 | 46.32 | 45.46 | 45.18 | 44.45 |
 | MDAA | MM-CTTA | 55.30 | 55.38 | 55.25 | 56.90 | 57.19 | 57.79 | 57.32 | 56.50 | 56.31 | 56.22 | 58.13 | 55.28 | 57.07 | 56.30 | 57.47 | 56.56 |
@@
 | SAR | 48.65 | 49.81 | 74.15 | 48.53 | 66.87 | 65.68 | 62.56 | 70.67 | 68.00 | 66.45 | 58.80 | 60.42 | 69.96 | 47.69 | 75.19 | 67.89 | 50.96 | 65.54 | 70.03 | 66.77 | 63.67 | 62.78 |
-| CoTTA | 50.21 | 47.72 | 72.16 | 44.96 | 58.80 | 55.53 | 58.28 | 61.35 | 61.52 | 59.73 | 50.15 | 53.92 | 63.15 | 49.48 | 65.16 | 56.71 | 43.97 | 55.47 | 49.68 | 54.79 | 60.14 | 55.85 |
 | EATA | 48.81 | 49.70 | 74.07 | 48.96 | 67.75 | 65.35 | 61.96 | 70.95 | 67.98 | 66.05 | 61.60 | 61.49 | 70.40 | 45.29 | 76.11 | 68.13 | 51.85 | 65.98 | 70.51 | 68.80 | 66.15 | 63.23 |
@@
 | READ | 51.18 | 53.62 | 73.88 | 54.73 | 68.67 | 67.95 | 67.07 | 70.14 | 68.84 | 67.62 | 62.68 | 64.90 | 68.24 | 59.46 | 71.43 | 68.16 | 52.60 | 66.35 | 66.18 | 62.85 | 63.35 | 64.28 |
+| READ+SRM | 46.18 | 45.39 | 55.52 | 44.25 | 56.45 | 44.43 | 55.42 | 59.41 | 49.61 | 57.63 | 54.45 | 53.65 | 50.92 | 50.60 | 62.39 | 52.52 | 45.93 | 59.55 | 52.69 | 55.92 | 56.06 | 52.81 |
 | MDAA | 55.04 | 55.91 | 73.64 | 55.89 | 63.78 | 73.12 | 63.54 | 67.62 | 74.90 | 67.00 | 65.60 | 67.21 | 75.44 | 64.89 | 72.69 | 76.94 | 65.45 | 69.21 | 76.95 | 70.94 | 71.16 | 67.95 |
@@
 | SAR | 47.97 | 48.73 | 72.21 | 48.46 | 66.39 | 66.75 | 63.69 | 70.28 | 68.00 | 66.33 | 59.24 | 59.85 | 70.17 | 46.42 | 75.52 | 67.59 | 51.12 | 65.23 | 70.24 | 68.31 | 66.18 | 62.79 |
-| CoTTA | 56.40 | 58.83 | 71.10 | 58.56 | 65.11 | 64.63 | 65.33 | 67.54 | 64.49 | 65.16 | 62.44 | 61.61 | 68.98 | 52.40 | 73.39 | 68.37 | 49.00 | 66.00 | 67.99 | 66.45 | 65.62 | 63.78 |
 | EATA | 48.76 | 49.76 | 73.89 | 48.86 | 67.75 | 65.20 | 61.89 | 70.67 | 68.28 | 66.16 | 61.64 | 61.44 | 70.33 | 45.57 | 75.96 | 68.30 | 51.88 | 65.87 | 70.51 | 68.59 | 65.98 | 63.20 |
@@
 | READ | 51.49 | 52.29 | 71.09 | 50.70 | 62.44 | 63.72 | 65.01 | 66.42 | 66.39 | 64.86 | 60.97 | 64.44 | 67.85 | 62.48 | 74.68 | 72.12 | 54.12 | 69.15 | 70.27 | 69.74 | 68.37 | 64.22 |
+| READ+SRM | 44.57 | 44.24 | 55.61 | 44.36 | 56.04 | 44.20 | 55.60 | 59.01 | 50.48 | 57.75 | 54.41 | 53.47 | 50.80 | 50.95 | 62.64 | 52.76 | 46.44 | 60.00 | 53.65 | 56.84 | 58.81 | 52.98 |
 | MDAA | 70.44 | 70.21 | 77.35 | 70.12 | 72.13 | 76.56 | 70.05 | 72.44 | 76.07 | 71.46 | 68.94 | 68.72 | 75.92 | 66.51 | 73.32 | 75.10 | 61.78 | 68.05 | 73.52 | 68.56 | 64.91 | 71.06 |
@@
 | SAR | 53.16 | 53.33 | 34.26 | 53.17 | 56.94 | 11.27 | 57.14 | 58.29 | 9.30 | 57.65 | 56.11 | 56.78 | 13.36 | 55.94 | 57.87 | 17.95 | 53.13 | 56.51 | 20.46 | 55.30 | 55.79 | 44.94 |
-| CoTTA | 52.24 | 52.26 | 10.66 | 49.15 | 49.22 | 1.68 | 46.15 | 46.38 | 1.34 | 44.84 | 44.66 | 44.13 | 0.65 | 43.57 | 43.23 | 4.83 | 35.16 | 34.85 | 0.56 | 36.10 | 35.34 | 32.24 |
 | EATA | 53.63 | 53.70 | 40.51 | 53.59 | 57.21 | 30.81 | 57.45 | 58.49 | 29.80 | 57.85 | 56.37 | 56.85 | 30.55 | 56.72 | 59.13 | 37.29 | 54.31 | 58.27 | 32.58 | 57.28 | 57.55 | 50.00 |
@@
 | READ | 53.78 | 53.91 | 39.83 | 54.17 | 57.81 | 26.00 | 58.14 | 59.42 | 21.63 | 59.03 | 57.38 | 58.29 | 22.79 | 57.71 | 59.32 | 26.07 | 55.46 | 58.36 | 18.29 | 57.36 | 57.63 | 48.21 |
+| READ+SRM | 42.34 | 41.70 | 22.88 | 41.73 | 45.39 | 10.93 | 45.62 | 46.75 | 11.23 | 46.28 | 45.00 | 44.82 | 10.92 | 42.59 | 47.43 | 19.22 | 40.47 | 46.43 | 12.57 | 45.02 | 45.39 | 35.94 |
 | MDAA | 55.09 | 55.31 | 38.60 | 55.31 | 56.89 | 34.83 | 57.20 | 57.69 | 34.65 | 57.35 | 56.47 | 56.27 | 34.28 | 56.17 | 58.05 | 36.86 | 55.33 | 57.00 | 35.53 | 56.28 | 57.35 | 50.60 |
@@
 | SAR | 51.90 | 51.78 | 25.90 | 51.30 | 54.73 | 5.11 | 54.82 | 56.52 | 7.82 | 56.31 | 54.63 | 55.49 | 13.44 | 55.12 | 57.59 | 15.07 | 53.57 | 56.92 | 15.04 | 56.53 | 57.30 | 43.19 |
-| CoTTA | 42.97 | 43.65 | 6.10 | 44.29 | 44.97 | 1.09 | 45.45 | 45.60 | 2.20 | 49.53 | 48.97 | 49.89 | 13.53 | 51.33 | 53.72 | 19.60 | 52.45 | 56.04 | 18.76 | 56.22 | 55.88 | 38.20 |
 | EATA | 53.77 | 53.65 | 40.39 | 53.61 | 57.17 | 30.49 | 57.36 | 58.59 | 30.16 | 57.82 | 56.21 | 56.64 | 31.22 | 56.77 | 59.24 | 37.44 | 54.30 | 58.18 | 33.01 | 57.46 | 57.53 | 50.05 |
@@
 | READ | 54.35 | 54.56 | 25.03 | 54.38 | 57.99 | 17.67 | 57.76 | 58.74 | 20.57 | 58.66 | 56.91 | 57.58 | 20.81 | 58.04 | 59.10 | 33.82 | 55.54 | 58.13 | 32.75 | 57.80 | 58.34 | 48.03 |
+| READ+SRM | 41.84 | 41.72 | 21.72 | 41.78 | 45.17 | 10.47 | 45.62 | 46.73 | 11.84 | 46.34 | 44.98 | 44.83 | 10.98 | 42.69 | 47.49 | 19.50 | 40.59 | 46.53 | 12.48 | 44.97 | 45.96 | 35.92 |
 | MDAA | 55.30 | 55.41 | 38.64 | 55.29 | 56.91 | 35.40 | 57.14 | 57.81 | 34.85 | 57.34 | 56.48 | 56.29 | 34.52 | 56.23 | 58.24 | 37.22 | 55.31 | 57.13 | 35.12 | 56.30 | 57.47 | 50.69 |
-4.3 Ablation Study
-In this section, we conduct three ablation studies on both video-corrupted Kinetics50-C dataset and audio-corrupted VGGSound-C dataset in severity level 5, with the batch size of 64.
-For simplicity, in the following section we use KS-video and VGG-audio to represent these two tasks.
+4. Experiments
+In this section, we evaluate MDAA under two challenging settings which is detailed in Sec.
+4.1.
+Sec.
+4.2 compares MDAA with other SOTA methods through extensive experiments and Sec.4.3 conducts ablation studies on MDAA’s components and computational efficiency.
+Implementation details are in Appendix B.
+4.1. Benchmarks and Settings
+To evaluate the model performance, we designed two tasks specifically for MM-CTTA.
+The first task named Progressive single-modality corruption, sequentially corrupts one modality while keeping the other clean, assessing the model’s resistance to catastrophic forgetting.
+It follows an online setting, where the model processes one sample at a time, leading to an extreme long learning phase.
+The second task, called interleaved modality corruption, continually alternates corruption between the two modalities.
+While most methods perform poorly in the online setting due to severe catastrophic forgetting, this task uses a batch size of 64 during test time to emphasize assessing the model’s ability to adapt to dynamic reliability biases.
+We utilize two datasets for both tasks: Kinetics50 (Kay et al., 2017) and VGGSound (Chen et al., 2020).
+Following (Yang et al., 2024), we construct their corrupted type named Kinetics50-C and VGGSound-C, each dataset includes 15 types of video corruptions and 6 audio corruptions at the largest severity (level 5).
+4.2. Performance Comparison
+In this section, we reproduce various TTA methods under the MM-CTTA setting.
+These include typical TTA methods (TENT (Wang et al., 2021), SAR (Niu et al., 2023)), CTTA methods (EATA (Niu et al., 2022)), and MM-TTA methods (MMTTA (Shin et al., 2022), READ (Xiong et al., 2024)).
+Additionally, we introduce an MM-CTTA variant, READ+SRM, which integrates READ with the stochastic restoration mechanism (SRM) from CoTTA (Wang et al., 2022).
+To ensure fairness, all methods use the same pre-trained model CAV-MAE (Gong et al., 2022b) as the multi-modal encoder.
+For methods that update BN layers, we instead update LN layers to align with the ViT structure of encoder.
+We also evaluate the source model, a baseline trained only on the source dataset and kept frozen during testing.
+The performances of various methods on progressive single-modality corruption are summarized in Tables 1, 2, and 3.
+Table 1 reports results on audio corruption, while Tables 2 and 3 focus on video corruption.
+Notably, most comparison methods collapse on VGGSound-C, performing worse than source models.
+In contrast, EATA achieves better results by limiting parameter updates, enabling more stable adaptation.
+MDAA outperforms prior methods by 3.00%-3.57% (audio) and 3.03%-6.22% (video) on Kinetics50-C, and by 0.94%-1.10% (audio) and 0.13%-0.18% (video) on VGGSound-C, on average.
+While MDAA exhibits suboptimal performance in scenarios where data corruption minimally impacts source model performance—attributable to its reduced parameter space for adaptation compared to baselines—it nevertheless demonstrates superior overall adaptability.
+MDAA consistently retains its advantage in later stages of the tasks, demonstrating superior robustness against catastrophic forgetting in MM-CTTA.
+The comparison results of interleaved modality corruption tasks are shown in Table 4 and 5.
+In this task, EATA, which is more good at memorization, is not dominant in the task of highlighting reliability bias.
+READ, which is specifically designed to address intra-modal reliability bias, demonstrates strong performance in this area.
+However, its effectiveness is limited to video corruption in MM-CTTA, as performance drops significantly during audio corruption.
+In contrast, MDAA is well-adapted to the corruption of different modalities, outperforming READ by 2.39%-6.84% and EATA by 0.60%-7.28% on average across Kinetics50-C and VGGSound-C.
+4.3. Ablation Study
+All studies in this section use the video-corrupted Kinetics50-C (KS-video) and audio-corrupted VGGSound-C (VGG-audio) datasets at severity level 5 with a batch size of 64.
 Component analysis.
 To verify the effectiveness of each MDAA component, we adopt an ablation study w.r.t three components as shown in Table 6.
-As observed, the model using only AC underperforms, with an average accuracy 0.5%-0.57% lower than READ on KS-video, and the model even collapses on VGG-audio, with an average accuracy of 0.47%-0.56%.
-This occurs because, while AC can prevent model forgetting, it cannot filter out unreliable samples, leading to issues with error accumulation and reliability bias.
-In contrast, the addition of DSM significantly improves performance, with gains of 1.73%-7.52% on KS-video and 33.74%-33.90% on VGG-audio.
-The introduction of SPS allows the model to learn from more possibly correct labels at the same time, thus further improving the performances on most tasks.
+As observed, due to the vulnerability of least square loss, the model using only ACs underperforms, with an average accuracy 0.5%-0.57% lower than READ (using cross-entropy loss) on KS-video.
+The model even collapses on VGG-audio, with an average accuracy of 0.47%-0.56%.
+The introduction of DLFM consistently improves model performance across all types of corruption, achieving the gains of 1.73%-7.52% on KS-video and 33.74%-33.90% on VGG-audio.
+Soft labeling introduces a small amount of uncertainty compared to hard labeling, it effectively alleviates the negative impact caused by incorrect pseudo labels during TTA.
+Although such uncertainty may lead to slightly lower performance in some tasks where corruptions in these tasks caused little effect on the model classification, soft labeling represents a favorable trade-off, particularly when dealing with highly complex datasets such as VGGSound.
+Furthermore, combining DLFM with soft labels yields the best overall performance.
+Reliable selection threshold.
+To examine the effect of the threshold on the DLFM, we plot the model’s performance with of from 1-5 to 1-1 in Figure 4(A-B).
+The performance of the model on both datasets exhibits an increasing and then decreasing trend.
+When is close to , the ACs are updated for nearly every sample, introducing more error.
+Conversely, when increases too much, the ACs do not learn from new inputs since no samples can pass through the DLFM, leading to a decline in performance.
+Therefore, value of should not be too small, otherwise the model will not be able to filter out the update samples with low execution.
+For both datasets in the experiment, a reasonable value of is around 1-3.
 | Method | KS-video | VGG-audio | ||
 | READ | 62.32 | 62.59 | 23.93 | 22.39 |
-| MDAA (ACs) | 61.82 | 62.02 | 0.47 | 0.56 |
-| MDAA (ACs+DSM) | 63.55 | 69.54 | 34.87 | 34.85 |
-| MDAA (ACs+DSM+SPS) | 65.43 | 69.30 | 35.82 | 35.77 |
-Reliable selection threshold.
-To examine the effect of the threshold on the DSM, we plot the model’s performance with of 0, 1-4, 5-4, 1-3, 2-3 and 5-3 in Fig.4(A-B).
-The performance of the model on both datasets exhibits an increasing and then decreasing trend.
-When is close to , the ACs are updated for nearly every sample, introducing more error.
-Conversely, when increases too much, the ACs do not learn from new inputs since no samples can pass through the DSM, leading to a decline in performance.
-Therefore, the we choose 1-3 a moderate valuefor both datasets as the threshold.
+| MDAA (ACs+Hard) | 61.82 | 62.02 | 0.47 | 0.56 |
+| MDAA (ACs+Soft) | 61.33 | 67.47 | 32.50 | 33.00 |
+| MDAA (ACs+DLFM+Hard) | 63.55 | 69.54 | 34.87 | 34.85 |
+| MDAA (ACs+DLFM+Soft) | 65.43 | 69.30 | 35.82 | 35.77 |
+| Method | KS-video | VGG-audio | ||
+| time(s) | memory(MB) | time(s) | memory(MB) | |
+| TENT* | 1085 | 37620 | 2571 | 37622 |
+| SAR | 2149 | 37620 | 4823 | 39220 |
+| EATA | 889 | 67286 | 2313 | 38300 |
+| MMTTA* | 2574 | 71732 | 5966 | 68840 |
+| READ | 1174 | 9176 | 2611 | 9178 |
+| MDAA | 1024 | 8784 | 1682 | 8750 |
 Soft label reconstruction.
-In this part we determine a suitable number to reconstruct the pseudo-labels.
-Given the sorted top distribution set ,we assign weights to each class in a decreasing manner using the formula .
-The results of the model are plotted in Fig.4(C-D) with chosen from to .
-While using soft labels can inevitably introduce error, there are instances where performance with one-hot labels may exceed that of the SPS (e.g., KS-video ).
-However, SPS remains beneficial as the dataset becomes larger and more complex.
-In fact, performance using SPS surpasses that of one-hot labels in most tasks, making SPS a worthwhile trade-off in MM-CTTA.
-Generally, the optimal number for reconstructions depends on the number of categories in the dataset, and a larger number of categories warrants a larger .
-However, more classes will be included in the reconstruction as increase, thereby introducing more error and reducing model performance.
-Therefore in SPS, we determine to use top and classes to reconstructed label in Kinetics50-C and VGGSound-C respectively.
-5 Conclusion
-In this paper, we analysed the factors that affect the model in the MM-CTTA task (i.e., error accumulation, catastrophic forgetting and reliability bias) and demonstrate that typical TTA methods are not suitable for the MM-CTTA task.
-To address the impact of these factors, we propose a new paradigm MDAA that introduce analytic learning to TTA for the first time.
-Instead of just adapting the model to the target domain, MDAA integrates the target domain into source domain, and thus prevent network from forgetting.
-With the help of DSM and SPS, model is able to dynamically and comprehensively process the information provided by each modality and use reliable samples to update.
-In the future, we will try to adapt this paradigm to more modalities to solve more challenging problems in real scenarios.
-Appendix A Proofs of Theorems
-In this section, we provide comprehensive proofs of Theorems 1 and 2 presented in the main paper.
-Note above equation as , derive for as
-Therefore .
-Proofs of Theorem 2. Similar to the proof of Theorem 1, we note the optimal formula as and derive it in terms of as
-Therefore .
-Appendix B Pseudo-code for MDAA
-In this appendix, we provide the pseudo-code for our MDAA in Algorithm 1.
-For the pre-trained model, we integrate an individual Analytic Classifier (AC) for each network block, using the source dataset to initialize the classifiers as well as the memory bank.
-During the inference and adaptation periods, the model reconstructs the output labels for each sample using the Soft Pseudo-label Strategy (SPS) and determines which ACs need to be updated through the Dynamic Selecting Mechanism (DSM).
-Appendix C Benchmarks,Backbone and Implementation Details
-C.1 Details about the Benchmarks
+Figure 4(C-D) shows the model results with ranging from to .
+While using soft labels inevitably introduces error, there are cases such as KS-video where one-hot labels may outperform soft labels.
+However, using soft pseudo-labels remains beneficial as datasets become larger and more complex.
+The choice of needs to be moderate, as larger datasets with more categories tend to require larger , while more error is likely to be introduced as increases.
+In this work, we determine to use top and classes for label reconstruction in Kinetics50-C and VGGSound-C respectively.
+Runtime and VRAM usage.
+To demonstrate MDAA’s practicality, we compare its runtime and VRAM usage against various methods in Table 7.
+MDAA achieves superior efficiency with lower memory overhead and faster execution, thanks to its unique design.
+Unlike other methods, MDAA eliminates backpropagation, reducing gradient computation and optimizer storage costs.
+The primary computational cost of MDAA lies in matrix inversion, with a time complexity of , where is the expanded feature dimension, fixed at 8000 in our experiments.
+5. Conclusion
+MM-CTTA shows great potential in real-world scenarios, as it enables pre-trained models to handle multi-modal inputs and adapt to continuously changing target domains.
+However, it also suffers from critical challenges such as reliability bias and catastrophic forgetting.
+To tackle these issues, we propose a new paradigm, MDAA, which introduces analytic learning into TTA.
+Instead of simply adapting the model to the target domain, MDAA integrates the target domain into source domain, and thus prevent network from forgetting.
+With the help of DLFM, model is able to dynamically and comprehensively process the information provided by each modality and use reliable samples to update.
+Moreover, compared to most SOTA TTA methods, MDAA achieves faster computational times and lower memory usage, highlighting its robust applicability in real-time deployment.
+In the future, we will try to adapt this method to more modalities to solve more challenging tasks.
+Acknowledgements.
+This research was supported by the National Natural Science Foundation of China (Grant No.
+62306117), the Guangzhou Basic and Applied Basic Research Foundation (2024A04J3681), and GJYC program of Guangzhou (2024D03J0005).References
+- (1)
+- Boudiaf et al.
+(2022) Malik Boudiaf, Romain Mueller, Ismail Ben Ayed, and Luca Bertinetto.
+2022.
+Parameter-free online test-time adaptation.
+In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition.
+8344–8353.
+- Cao et al.
+(2023) Haozhi Cao, Yuecong Xu, Jianfei Yang, Pengyu Yin, Shenghai Yuan, and Lihua Xie.
+2023.
+Multi-modal continual test-time adaptation for 3d semantic segmentation.
+In Proceedings of the IEEE/CVF International Conference on Computer Vision.
+18809–18819.
+- Chen et al.
+(2019) Chaoqi Chen, Weiping Xie, Wenbing Huang, Yu Rong, Xinghao Ding, Yue Huang, Tingyang Xu, and Junzhou Huang.
+2019.
+Progressive feature alignment for unsupervised domain adaptation.
+In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition.
+627–636.
+- Chen et al.
+(2020) Honglie Chen, Weidi Xie, Andrea Vedaldi, and Andrew Zisserman.
+2020.
+Vggsound: A large-scale audio-visual dataset.
+In ICASSP 2020-2020 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP).
+IEEE, 721–725.
+- Dosovitskiy et al.
+(2020) Alexey Dosovitskiy, Lucas Beyer, Alexander Kolesnikov, Dirk Weissenborn, Xiaohua Zhai, Thomas Unterthiner, Mostafa Dehghani, Matthias Minderer, Georg Heigold, Sylvain Gelly, et al.
+2020.
+An image is worth 16x16 words: Transformers for image recognition at scale.
+In International Conference on Learning Representations.
+- Fang et al. (2024) Di Fang, Yinan Zhu, Runze Fang, Cen Chen, Ziqian Zeng, and Huiping Zhuang. 2024. AIR: Analytic Imbalance Rectifier for Continual Learning. arXiv preprint arXiv:2408.10349 (2024).
+- Gao et al.
+(2023) Jin Gao, Jialing Zhang, Xihui Liu, Trevor Darrell, Evan Shelhamer, and Dequan Wang.
+2023.
+Back to the source: Diffusion-driven adaptation to test-time corruption.
+In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition.
+11786–11796.
+- Gong et al.
+(2022a) Taesik Gong, Jongheon Jeong, Taewon Kim, Yewon Kim, Jinwoo Shin, and Sung-Ju Lee.
+2022a.
+Note: Robust continual test-time adaptation against temporal correlation.
+Advances in Neural Information Processing Systems 35 (2022), 27253–27266.
+- Gong et al.
+(2022b) Yuan Gong, Andrew Rouditchenko, Alexander H Liu, David Harwath, Leonid Karlinsky, Hilde Kuehne, and James Glass.
+2022b.
+Contrastive audio-visual masked autoencoder.
+arXiv preprint arXiv:2210.07839 (2022).
+- Guo and Lyu (2004) Ping Guo and Michael R Lyu.
+2004.
+A pseudoinverse learning algorithm for feedforward neural networks with stacked generalization applications to software reliability growth data.
+Neurocomputing 56 (2004), 101–121.
+- Guo and Jin (2025) Zirun Guo and Tao Jin. 2025. Smoothing the shift: Towards stable test-time adaptation under complex multimodal noises. arXiv preprint arXiv:2503.02616 (2025).
+- Hinton et al. (2015) Geoffrey Hinton, Oriol Vinyals, and Jeff Dean. 2015. Distilling the Knowledge in a Neural Network. stat 1050 (2015), 9.
+- Kay et al.
+(2017) Will Kay, Joao Carreira, Karen Simonyan, Brian Zhang, Chloe Hillier, Sudheendra Vijayanarasimhan, Fabio Viola, Tim Green, Trevor Back, Paul Natsev, et al.
+2017.
+The kinetics human action video dataset.
+arXiv preprint arXiv:1705.06950 (2017).
+- Lei and Pernkopf (2024) Jixiang Lei and Franz Pernkopf. 2024. Two-Level Test-Time Adaptation in Multimodal Learning. In ICML 2024 Workshop on Foundation Models in the Wild.
+- Liang et al. (2024) Jian Liang, Ran He, and Tieniu Tan. 2024. A comprehensive survey on test-time adaptation under distribution shifts. International Journal of Computer Vision (2024), 1–34.
+- McCloskey and Cohen (1989) Michael McCloskey and Neal J Cohen.
+1989.
+Catastrophic interference in connectionist networks: The sequential learning problem.
+In Psychology of learning and motivation.
+Vol.
+24.
+Elsevier, 109–165.
+- Müller et al. (2019) Rafael Müller, Simon Kornblith, and Geoffrey E Hinton. 2019. When does label smoothing help? Advances in neural information processing systems 32 (2019).
+- Niu et al.
+(2022) Shuaicheng Niu, Jiaxiang Wu, Yifan Zhang, Yaofo Chen, Shijian Zheng, Peilin Zhao, and Mingkui Tan.
+2022.
+Efficient test-time model adaptation without forgetting.
+In International conference on machine learning.
+PMLR, 16888–16905.
+- Niu et al.
+(2023) Shuaicheng Niu, Jiaxiang Wu, Yifan Zhang, Zhiquan Wen, Yaofo Chen, Peilin Zhao, and Mingkui Tan.
+2023.
+Towards Stable Test-time Adaptation in Dynamic Wild World.
+In The Eleventh International Conference on Learning Representations.
+- Radford et al.
+(2021) Alec Radford, Jong Wook Kim, Chris Hallacy, Aditya Ramesh, Gabriel Goh, Sandhini Agarwal, Girish Sastry, Amanda Askell, Pamela Mishkin, Jack Clark, et al.
+2021.
+Learning transferable visual models from natural language supervision.
+In International conference on machine learning.
+PMLR, 8748–8763.
+- Shin et al.
+(2022) Inkyu Shin, Yi-Hsuan Tsai, Bingbing Zhuang, Samuel Schulter, Buyu Liu, Sparsh Garg, In So Kweon, and Kuk-Jin Yoon.
+2022.
+Mm-tta: multi-modal test-time adaptation for 3d semantic segmentation.
+In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition.
+16928–16937.
+- Vaswani (2017) A Vaswani. 2017. Attention is all you need. Advances in Neural Information Processing Systems (2017).
+- Wang et al.
+(2021) Dequan Wang, Evan Shelhamer, Shaoteng Liu, Bruno Olshausen, and Trevor Darrell.
+2021.
+Tent: Fully Test-Time Adaptation by Entropy Minimization.
+In International Conference on Learning Representations.
+- Wang et al.
+(2022) Qin Wang, Olga Fink, Luc Van Gool, and Dengxin Dai.
+2022.
+Continual test-time domain adaptation.
+In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition.
+7201–7211.
+- Wang et al.
+(2020) Weiyao Wang, Du Tran, and Matt Feiszli.
+2020.
+What makes training multi-modal classification networks hard?.
+In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition.
+12695–12705.
+- Xiong et al.
+(2024) Baochen Xiong, Xiaoshan Yang, Yaguang Song, Yaowei Wang, and Changsheng Xu.
+2024.
+Modality-Collaborative Test-Time Adaptation for Action Recognition.
+In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition.
+26732–26741.
+- Yang et al.
+(2024) Mouxing Yang, Yunfan Li, Changqing Zhang, Peng Hu, and Xi Peng.
+2024.
+Test-time Adaptation against Multi-modal Reliability Bias.
+In The Twelfth International Conference on Learning Representations.
+- Yang and Hu (2008) Shuang-Hong Yang and Bao-Gang Hu.
+2008.
+A stagewise least square loss function for classification.
+In Proceedings of the 2008 SIAM International Conference on Data Mining.
+SIAM, 120–131.
+- Zhang et al. (2015) Xu Zhang, Felix Xinnan Yu, Shih-Fu Chang, and Shengjin Wang. 2015. Deep transfer network: Unsupervised domain adaptation. arXiv preprint arXiv:1503.00591 (2015).
+- Zhang et al.
+(2023) Yifan Zhang, Xue Wang, Kexin Jin, Kun Yuan, Zhang Zhang, Liang Wang, Rong Jin, and Tieniu Tan.
+2023.
+Adanpc: Exploring non-parametric classifier for test-time adaptation.
+In International Conference on Machine Learning.
+PMLR, 41647–41676.
+- Zhao et al.
+(2025) Yusheng Zhao, Junyu Luo, Xiao Luo, Jinsheng Huang, Jingyang Yuan, Zhiping Xiao, and Ming Zhang.
+2025.
+Attention Bootstrapping for Multi-Modal Test-Time Adaptation.
+In Proceedings of the AAAI Conference on Artificial Intelligence, Vol.
+39.
+22849–22857.
+- Zhou et al.
+(2023) Zhi Zhou, Lan-Zhe Guo, Lin-Han Jia, Dingchu Zhang, and Yu-Feng Li.
+2023.
+ODS: Test-Time Adaptation in the Presence of Open-World Data Shift.
+In Proceedings of the 40th International Conference on Machine Learning (Proceedings of Machine Learning Research, Vol.
+202), Andreas Krause, Emma Brunskill, Kyunghyun Cho, Barbara Engelhardt, Sivan Sabato, and Jonathan Scarlett (Eds.).
+PMLR, 42574–42588.
+- Zhuang et al.
+(2021) Huiping Zhuang, Zhiping Lin, and Kar-Ann Toh.
+2021.
+Blockwise recursive Moore–Penrose inverse for network learning.
+IEEE Transactions on Systems, Man, and Cybernetics: Systems 52, 5 (2021), 3237–3250.
+- Zhuang et al.
+(2023) Huiping Zhuang, Zhenyu Weng, Run He, Zhiping Lin, and Ziqian Zeng.
+2023.
+Gkeal: Gaussian kernel embedded analytic learning for few-shot class incremental task.
+In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition.
+7746–7755.
+- Zhuang et al.
+(2022) Huiping Zhuang, Zhenyu Weng, Hongxin Wei, Renchunzi Xie, Kar-Ann Toh, and Zhiping Lin.
+2022.
+ACIL: Analytic class-incremental learning with absolute memorization and privacy protection.
+Advances in Neural Information Processing Systems 35 (2022), 11602–11614.
+Appendix A Pseudo-code for MDAA
+The pseudo-code of MDAA is shown in Algorithm 1.
+For the pre-trained models, we integrate an individual AC for each network block, using the source dataset to initialize the classifiers as well as the memory bank.
+During the inference and adaptation periods, the model reconstructs the output labels for each sample using the soft pseudo-label and determines which ACs need to be updated through the DLFM.
+Appendix B Benchmarks,Backbone and Implementation Details
+B.1. Details about the Benchmarks
 ALL experiments are conducted on the two popular multi-modal datasets Kinetics (Kay et al., 2017) and VGGSound (Chen et al., 2020).
-Yang et al.
-(2024) further provides their corrupted visual and audio modality for TTA tasks.
+(Yang et al., 2024) further provides their corrupted visual and audio modality for TTA tasks.
 Kinetics50 is a subset of the Kinetics dataset (Kay et al., 2017), consisting of 50 randomly selected classes (Yang et al., 2024).
@@
 Compared to Kinetics50, VGGSound includes a wider range of categories, introducing additional complexity to the classification task.
-Both datasets’ visual and audio modalities were extracted following the method described in Gong et al.
-(2022b).
+Both datasets’ visual and audio modalities were extracted following the method described in (Gong et al., 2022b).
 To systematically explore the distributional shifts across modalities, various corruption types were applied to both visual and audio components.
-Following Yang et al.
-(2024), 15 corruption types were introduced for the visual modality, each with five levels of severity for comprehensive evaluation.
-These corruptions include ”gaussian noise”, ”shot noise”, ”impulse noise”, ”defocus blur”, ”glass blur”, ”motion blur”, ”zoom blur”, ”snow”, ”frost”, ”fog”, ”brightness”, ”contrast”, ”elastic transform”, ”pixelate”, and ”jpeg compression”.
-Similarly, the audio modality was subjected to 6 different corruptions: ”gaussian noise”, ”traffic noise”, ”crowd noise”, ”rain”, ”thunder”, and ”wind”.
+Following (Yang et al., 2024), 15 corruption types were introduced for the visual modality, each with five levels of severity for comprehensive evaluation.
+These corruptions include “gaussian nois”, “shot noise”, “impulse noise”, “defocus blur”, “glass blur”, “motion blur”, “zoom blur”, “snow”, “frost”, “fog”, “brightness”, “contrast”, “elastic transform”, “pixelate”, and “jpeg compression”.
+Similarly, the audio modality was subjected to 6 different corruptions: “gaussian noise”, “traffic noise”, “crowd noise”, “rain”, “thunder”, and “wind”.
 The corrupted versions of these benchmarks are referred to as Kinetics50-C and VGGSound-C, respectively.
-Visualizations of sample corrupted video frames and audio spectrograms are provided in Fig.5 and Fig.6.
-C.2 CAV-MAE Backbone
-CAV-MAE is employed as the pre-trained model for MM-CTTA in this paper.
+Visualizations of sample corrupted video frames and audio spectrograms are provided in Figure 5 and Figure 6.
+B.2. CAV-MAE Backbone
+CAV-MAE is employed as the pre-trained models for MM-CTTA in this paper.
 Its architecture consists of 11 Transformer blocks (known as feature encoder networks) dedicated to modality-specific feature extraction, followed by an additional Transformer block (known as fusion network) responsible for cross modal fusion.
@@
 In the case of the audio input, the original 10-second audio waveform is transformed into a 2D spectrogram before being processed by the Transformer encoder for the audio modality (Gong et al., 2022a).
-C.3 Implementation Details
+| KS-video | VGG-audio | |||
+| 1e-3 | 2.10 | 2.08 | 0.51 | 0.47 |
+| 1e-2 | 2.03 | 2.12 | 0.31 | 0.39 |
+| 1e-1 | 64.87 | 69.03 | 0.32 | 0.37 |
+| 1e0 | 65.43 | 69.30 | 0.36 | 0.40 |
+| 1e1 | 65.27 | 69.23 | 35.82 | 35.77 |
+| 1e2 | 65.29 | 68.95 | 35.78 | 35.76 |
+| 1e3 | 59.32 | 59.94 | 30.36 | 30.52 |
+| Method | Type | Kinetics50-C | VGGSound-C | ||||||||||||
+| Gauss. | Traff. | Crowd | Rain | Thund. | Wind | Avg. | Gauss. | Traff. | Crowd | Rain | Thund. | Wind | Avg. | ||
+| Source | - | 73.97 | 65.21 | 67.79 | 70.27 | 67.98 | 70.45 | 69.28 | 37.32 | 21.24 | 16.89 | 21.82 | 27.37 | 25.66 | 25.05 |
+| TENT* | TTA | 74.44 | 68.04 | 71.30 | 70.25 | 72.53 | 70.35 | 71.15 | 10.76 | 1.15 | 0.40 | 0.32 | 0.51 | 0.31 | 2.24 |
+| SAR | TTA | 73.88 | 65.68 | 68.00 | 70.91 | 69.07 | 70.45 | 69.66 | 37.39 | 8.57 | 7.03 | 12.58 | 10.77 | 13.71 | 15.01 |
+| EATA | CTTA | 73.95 | 65.26 | 68.03 | 70.45 | 68.17 | 70.48 | 69.39 | 40.49 | 31.07 | 31.98 | 31.40 | 38.26 | 33.84 | 34.51 |
+| MMTTA* | MM-TTA | 69.32 | 69.01 | 69.00 | 69.07 | 68.96 | 68.95 | 69.05 | 14.40 | 1.92 | 0.84 | 0.36 | 0.47 | 0.31 | 3.05 |
+| READ | MM-TTA | 74.74 | 68.88 | 70.43 | 70.69 | 72.31 | 69.73 | 71.13 | 40.51 | 25.39 | 20.38 | 20.06 | 21.14 | 16.07 | 23.93 |
+| MDAA | MM-CTTA | 73.33 | 71.99 | 73.36 | 73.26 | 74.24 | 73.76 | 73.32 | 38.57 | 34.57 | 34.37 | 34.30 | 37.40 | 35.70 | 35.82 |
+| Source | - | 74.00 | 65.20 | 67.92 | 70.24 | 68.01 | 70.43 | 69.30 | 37.31 | 21.25 | 16.89 | 21.81 | 27.37 | 25.66 | 25.05 |
+| TENT* | TTA | 73.31 | 70.00 | 71.57 | 70.30 | 68.87 | 71.09 | 70.86 | 0.37 | 0.30 | 0.30 | 0.31 | 0.97 | 3.73 | 1.00 |
+| SAR | TTA | 73.29 | 66.79 | 68.14 | 70.71 | 67.97 | 70.36 | 69.55 | 32.02 | 9.92 | 7.04 | 9.49 | 11.41 | 16.03 | 14.32 |
+| EATA | CTTA | 73.96 | 65.30 | 68.15 | 70.36 | 68.18 | 70.40 | 69.39 | 40.65 | 32.32 | 30.65 | 32.23 | 38.16 | 32.39 | 34.40 |
+| MMTTA* | MM-TTA | 68.92 | 69.47 | 70.50 | 69.59 | 69.62 | 69.63 | 69.62 | 0.35 | 0.32 | 0.30 | 0.23 | 0.67 | 3.25 | 0.85 |
+| READ | MM-TTA | 73.11 | 70.16 | 69.68 | 70.90 | 72.07 | 70.73 | 71.11 | 24.23 | 16.23 | 16.84 | 17.65 | 29.12 | 30.30 | 22.39 |
+| MDAA | MM-CTTA | 72.55 | 73.67 | 72.98 | 73.04 | 73.18 | 71.55 | 72.83 | 38.59 | 35.35 | 34.50 | 34.31 | 36.94 | 34.91 | 35.77 |
+B.3. Implementation Details
 In the final step of Algorithm 1, we determine the hyperparameters for MDAA.
 The expansion layer dimension, denoted as , theoretically benefits from larger values.
-However, an excessively large dimension may introduce a significant number of parameters, increasing the computational load.
+However, an excessively large dimension may introduce a significant number of parameters, occupying considerable computational resources.
 Given the constraints of our available GPU resources, we set the dimension of to 8000.
-The necessity of the parameter in 3.2 has been established in Zhuang et al.
-(2022).
+The necessity of the parameter in Eq.
+3.2 has been established in (Zhuang et al., 2022).
 The model demonstrates stable performance over a wide range of values, indicating that as long as is within a reasonable range, its impact on the model’s performance remains minimal.
-After conducting a sweep over five orders of magnitude (), we set to 1 for Kinetics-50 and 10 for VGGSound.
-As discussed in Sections 4.2 and 4.3, the threshold in DSM is fixed at 0.001 for both datasets, while the parameter in SPS is set to 2 for Kinetics-50 and 7 for VGGSound.
+Table 8 shows the results under a sweep over six orders of magnitude (i.e., ) using the same setting in the ablation study.
+The model performs robustly with a sufficient regularization, while collapses when the value of turns to be too extreme (larger than and smaller than ).
+In this paper, we set to 1 for Kinetics-50 and 10 for VGGSound.
+As discussed in Sections 4.2 and 4.3, the threshold in DLFM is fixed at 0.001 for both datasets, while the parameter for reconstructing soft labels is set to 2 for Kinetics-50 and 7 for VGGSound.
 All experiments were conducted on an RTX3090 GPU, with results averaged over three runs.
-Appendix D More Experiment Results
-In this appendix we provide four more experiments for reference where Appendix .D.1 and D.2 are the supplementary comparative studies while Appendix .D.3 and D.4 are the further discussion on hyperparameters applied in MDAA.
-D.1 Performance Comparison on Multi-modal Corruption
-We follow the experimental setup in (Yang et al., 2024) to examine whether our model remains superior to others when the adaptation task is not continual.
-In this section, the corrupted target domains are treated as independent, and the results for corrupted audio and video modalities are presented in Tables 7, 8, and 9.
-The results show that the SOTA performance of MDAA is not only due to its ability to combat catastrophic forgetting but also its strong capacity to handle MM-TTA tasks effectively.
+Appendix C More Experiment Results
+In this appendix, we provide four more experiments for reference where Appendix C.1 and C.2 are the supplementary comparative studies while Appendix C.3 is the further discussion on hyperparameters applied in MDAA.
+| Method | Type | Gauss. | Shot. | Impul. | Defoc. | Glass. | Motion. | Zoom. | Snow | Frost | Fog | Bright. | Cont. | Elastic. | Pixel. | Jpeg | Avg. |
+| Source | - | 48.67 | 49.81 | 49.01 | 67.77 | 61.88 | 70.95 | 66.19 | 61.39 | 61.42 | 45.35 | 75.94 | 51.86 | 65.81 | 68.77 | 66.10 | 60.73 |
+| TENT* | TTA | 48.53 | 48.65 | 46.06 | 62.91 | 62.15 | 65.77 | 63.83 | 54.78 | 54.64 | 33.79 | 36.79 | 19.86 | 11.75 | 3.82 | 3.38 | 41.11 |
+| SAR | TTA | 48.48 | 49.87 | 48.71 | 66.92 | 62.56 | 70.58 | 66.77 | 59.25 | 60.50 | 47.19 | 75.34 | 50.77 | 65.23 | 66.91 | 64.01 | 60.21 |
+| EATA | CTTA | 48.76 | 49.84 | 49.03 | 67.78 | 62.02 | 70.92 | 66.20 | 61.55 | 61.53 | 45.38 | 75.97 | 51.90 | 65.91 | 68.74 | 66.09 | 60.77 |
+| MMTTA* | MM-TTA | 48.74 | 49.05 | 48.88 | 49.12 | 48.94 | 48.88 | 48.86 | 48.92 | 48.92 | 48.86 | 49.03 | 48.79 | 48.88 | 48.80 | 48.92 | 48.91 |
+| READ | MM-TTA | 51.02 | 53.54 | 54.24 | 68.16 | 66.36 | 68.95 | 67.76 | 62.85 | 64.72 | 59.62 | 70.99 | 53.02 | 67.03 | 63.71 | 62.78 | 62.32 |
+| MDAA | MM-CTTA | 55.84 | 55.79 | 55.50 | 64.29 | 63.71 | 68.04 | 68.10 | 66.06 | 68.01 | 65.71 | 72.97 | 66.36 | 70.04 | 70.51 | 70.53 | 65.43 |
+| Source | - | 48.68 | 49.79 | 48.97 | 67.69 | 61.82 | 70.93 | 66.18 | 61.39 | 61.37 | 45.29 | 75.99 | 51.89 | 65.78 | 68.71 | 66.14 | 60.71 |
+| TENT* | TTA | 41.93 | 47.19 | 49.22 | 64.80 | 64.33 | 68.05 | 63.86 | 61.42 | 62.12 | 51.65 | 73.92 | 50.80 | 67.47 | 69.03 | 66.80 | 60.17 |
+| SAR | TTA | 48.08 | 49.07 | 48.83 | 66.25 | 63.02 | 70.39 | 66.27 | 59.53 | 60.31 | 46.28 | 75.62 | 51.23 | 65.47 | 68.47 | 66.14 | 60.33 |
+| EATA | CTTA | 48.67 | 49.80 | 49.16 | 67.68 | 62.02 | 70.92 | 66.13 | 61.57 | 61.46 | 45.17 | 76.02 | 51.95 | 65.90 | 68.71 | 66.22 | 60.76 |
+| MMTTA* | MM-TTA | 49.82 | 49.90 | 49.85 | 49.94 | 49.86 | 49.92 | 49.89 | 49.97 | 49.94 | 49.79 | 50.54 | 50.01 | 54.33 | 56.20 | 59.99 | 51.33 |
+| READ | MM-TTA | 52.00 | 52.28 | 51.02 | 62.43 | 64.47 | 66.04 | 65.38 | 61.39 | 65.10 | 62.96 | 75.14 | 53.59 | 68.89 | 69.95 | 68.20 | 62.59 |
+| MDAA | MM-CTTA | 70.55 | 70.43 | 70.08 | 71.81 | 70.30 | 72.19 | 70.88 | 69.38 | 69.35 | 67.48 | 73.16 | 62.07 | 68.63 | 67.98 | 65.21 | 69.30 |
+| Method | Type | Gauss. | Shot. | Impul. | Defoc. | Glass. | Motion. | Zoom. | Snow | Frost | Fog | Bright. | Cont. | Elastic. | Pixel. | Jpeg | Avg. |
+| Source | - | 53.05 | 52.91 | 52.98 | 57.19 | 57.37 | 58.37 | 57.45 | 56.37 | 56.47 | 55.41 | 59.19 | 53.75 | 57.19 | 56.44 | 57.33 | 56.10 |
+| TENT* | TTA | 53.27 | 52.76 | 52.00 | 54.58 | 54.35 | 55.04 | 54.86 | 52.59 | 52.81 | 53.11 | 53.50 | 50.80 | 53.03 | 52.31 | 52.14 | 53.14 |
+| SAR | TTA | 53.14 | 53.29 | 53.21 | 56.95 | 56.91 | 58.41 | 57.47 | 56.10 | 56.78 | 56.34 | 58.35 | 53.98 | 57.27 | 55.93 | 56.48 | 56.04 |
+| EATA | CTTA | 53.74 | 53.67 | 53.68 | 57.20 | 57.26 | 58.53 | 57.93 | 56.38 | 56.67 | 56.23 | 59.04 | 53.63 | 58.19 | 57.36 | 57.48 | 56.47 |
+| READ | MM-TTA | 53.77 | 54.26 | 54.28 | 58.04 | 58.00 | 59.09 | 58.84 | 57.43 | 58.18 | 58.12 | 59.38 | 55.99 | 58.30 | 57.51 | 57.91 | 57.27 |
+| MDAA | MM-CTTA | 55.63 | 55.91 | 55.88 | 57.50 | 57.77 | 58.27 | 57.84 | 57.05 | 56.72 | 56.77 | 58.67 | 55.78 | 57.55 | 56.87 | 57.84 | 57.07 |
+| Source | - | 53.01 | 52.88 | 52.97 | 57.20 | 57.42 | 58.41 | 57.49 | 56.37 | 56.49 | 55.43 | 59.16 | 53.74 | 57.19 | 56.42 | 57.29 | 56.10 |
+| TENT* | TTA | 50.48 | 50.50 | 50.50 | 53.09 | 53.19 | 53.70 | 53.18 | 52.31 | 53.47 | 53.57 | 55.40 | 52.22 | 56.23 | 56.74 | 56.99 | 53.44 |
+| SAR | TTA | 53.02 | 52.95 | 52.71 | 56.78 | 56.63 | 58.18 | 57.38 | 55.97 | 56.69 | 56.45 | 58.46 | 54.03 | 57.41 | 56.53 | 57.16 | 56.02 |
+| EATA | CTTA | 53.69 | 53.66 | 53.61 | 57.26 | 57.34 | 58.46 | 57.84 | 56.35 | 56.82 | 56.72 | 59.21 | 54.10 | 58.28 | 57.39 | 57.54 | 56.55 |
+| READ | MM-TTA | 55.33 | 55.34 | 54.93 | 58.16 | 57.88 | 58.82 | 58.50 | 57.24 | 57.94 | 57.86 | 59.65 | 55.67 | 58.79 | 57.87 | 58.10 | 57.47 |
+| MDAA | MM-CTTA | 55.85 | 55.87 | 55.94 | 57.40 | 57.85 | 58.25 | 57.95 | 57.02 | 56.89 | 56.68 | 58.66 | 55.84 | 57.58 | 56.92 | 58.00 | 57.11 |
+C.1. Performance Comparison on Multi-modal Corruption
+We follow the experimental setup in (Yang et al., 2024) to examine whether our model remains superior to others when the adaptation task is not continuous.
+In this section, the corrupted target domains are independent, and the results for corrupted audio and video modalities are presented in Tables 9, 10 and 11.
+The results show that the SOTA performance of MDAA is due not only to its ability to combat catastrophic forgetting, but also to its strong capacity to handle MM-TTA tasks effectively.
 | Method | Type | Kinetics50-C | VGGSound-C | ||||||||||||
@@
 | MDAA | MM-CTTA | 54.89 | 55.25 | 55.32 | 63.89 | 62.49 | 67.26 | 65.86 | 64.32 | 65.31 | 61.86 | 73.20 | 61.60 | 67.83 | 69.22 | 68.69 | 63.80 |
-D.2 Performance Comparison on Single-modality Continual Corruption
+C.2. Performance Comparison on Single-modality Continual Corruption
 In this section, we compare the performance of each method under single-modality corruption.
 This task is similar to the progressive single-modality corruption task described in the main text, but here we use a batch size of 64.
-| Method | Type | Kinetics50-C | VGGSound-C | ||||||||||||
-| Gauss. | Traff. | Crowd | Rain | Thund. | Wind | Avg. | Gauss. | Traff. | Crowd | Rain | Thund. | Wind | Avg. | ||
-| Source | - | 73.97 | 65.21 | 67.79 | 70.27 | 67.98 | 70.45 | 69.28 | 37.32 | 21.24 | 16.89 | 21.82 | 27.37 | 25.66 | 25.05 |
-| TENT* | TTA | 74.44 | 68.04 | 71.30 | 70.25 | 72.53 | 70.35 | 71.15 | 10.76 | 1.15 | 0.40 | 0.32 | 0.51 | 0.31 | 2.24 |
-| SAR | TTA | 73.88 | 65.68 | 68.00 | 70.91 | 69.07 | 70.45 | 69.66 | 37.39 | 8.57 | 7.03 | 12.58 | 10.77 | 13.71 | 15.01 |
-| CoTTA | CTTA | 73.89 | 66.84 | 68.08 | 67.53 | 71.10 | 69.33 | 69.46 | 39.70 | 34.88 | 35.54 | 33.67 | 39.42 | 33.83 | 36.17 |
-| EATA | CTTA | 73.95 | 65.26 | 68.03 | 70.45 | 68.17 | 70.48 | 69.39 | 40.49 | 31.07 | 31.98 | 31.40 | 38.26 | 33.84 | 34.51 |
-| MMTTA* | MM-TTA | 69.32 | 69.01 | 69.00 | 69.07 | 68.96 | 68.95 | 69.05 | 14.40 | 1.92 | 0.84 | 0.36 | 0.47 | 0.31 | 3.05 |
-| READ | MM-TTA | 74.74 | 68.88 | 70.43 | 70.69 | 72.31 | 69.73 | 71.13 | 40.51 | 25.39 | 20.38 | 20.06 | 21.14 | 16.07 | 23.93 |
-| MDAA | MM-CTTA | 73.33 | 71.99 | 73.36 | 73.26 | 74.24 | 73.76 | 73.32 | 38.57 | 34.57 | 34.37 | 34.30 | 37.40 | 35.70 | 35.82 |
-| Source | - | 74.00 | 65.20 | 67.92 | 70.24 | 68.01 | 70.43 | 69.30 | 37.31 | 21.25 | 16.89 | 21.81 | 27.37 | 25.66 | 25.05 |
-| TENT* | TTA | 73.31 | 70.00 | 71.57 | 70.30 | 68.87 | 71.09 | 70.86 | 0.37 | 0.30 | 0.30 | 0.31 | 0.97 | 3.73 | 1.00 |
-| SAR | TTA | 73.29 | 66.79 | 68.14 | 70.71 | 67.97 | 70.36 | 69.55 | 32.02 | 9.92 | 7.04 | 9.49 | 11.41 | 16.03 | 14.32 |
-| CoTTA | CTTA | 69.88 | 67.67 | 67.84 | 68.54 | 69.63 | 70.76 | 69.05 | 25.06 | 25.08 | 28.01 | 33.22 | 37.90 | 29.67 | 29.82 |
-| EATA | CTTA | 73.96 | 65.30 | 68.15 | 70.36 | 68.18 | 70.40 | 69.39 | 40.65 | 32.32 | 30.65 | 32.23 | 38.16 | 32.39 | 34.40 |
-| MMTTA* | MM-TTA | 68.92 | 69.47 | 70.50 | 69.59 | 69.62 | 69.63 | 69.62 | 0.35 | 0.32 | 0.30 | 0.23 | 0.67 | 3.25 | 0.85 |
-| READ | MM-TTA | 73.11 | 70.16 | 69.68 | 70.90 | 72.07 | 70.73 | 71.11 | 24.23 | 16.23 | 16.84 | 17.65 | 29.12 | 30.30 | 22.39 |
-| MDAA | MM-CTTA | 72.55 | 73.67 | 72.98 | 73.04 | 73.18 | 71.55 | 72.83 | 38.59 | 35.35 | 34.50 | 34.31 | 36.94 | 34.91 | 35.77 |
-| Method | Type | Gauss. | Shot. | Impul. | Defoc. | Glass. | Motion. | Zoom. | Snow | Frost | Fog | Bright. | Cont. | Elastic. | Pixel. | Jpeg | Avg. |
-| Source | - | 48.67 | 49.81 | 49.01 | 67.77 | 61.88 | 70.95 | 66.19 | 61.39 | 61.42 | 45.35 | 75.94 | 51.86 | 65.81 | 68.77 | 66.10 | 60.73 |
-| TENT* | TTA | 48.53 | 48.65 | 46.06 | 62.91 | 62.15 | 65.77 | 63.83 | 54.78 | 54.64 | 33.79 | 36.79 | 19.86 | 11.75 | 3.82 | 3.38 | 41.11 |
-| SAR | TTA | 48.48 | 49.87 | 48.71 | 66.92 | 62.56 | 70.58 | 66.77 | 59.25 | 60.50 | 47.19 | 75.34 | 50.77 | 65.23 | 66.91 | 64.01 | 60.21 |
-| CoTTA | CTTA | 49.17 | 46.65 | 43.69 | 61.82 | 60.00 | 63.54 | 62.34 | 50.60 | 52.13 | 53.35 | 60.25 | 50.86 | 60.43 | 58.28 | 62.61 | 55.71 |
-| EATA | CTTA | 48.76 | 49.84 | 49.03 | 67.78 | 62.02 | 70.92 | 66.20 | 61.55 | 61.53 | 45.38 | 75.97 | 51.90 | 65.91 | 68.74 | 66.09 | 60.77 |
-| MMTTA* | MM-TTA | 48.74 | 49.05 | 48.88 | 49.12 | 48.94 | 48.88 | 48.86 | 48.92 | 48.92 | 48.86 | 49.03 | 48.79 | 48.88 | 48.80 | 48.92 | 48.91 |
-| READ | MM-TTA | 51.02 | 53.54 | 54.24 | 68.16 | 66.36 | 68.95 | 67.76 | 62.85 | 64.72 | 59.62 | 70.99 | 53.02 | 67.03 | 63.71 | 62.78 | 62.32 |
-| MDAA | MM-CTTA | 55.84 | 55.79 | 55.50 | 64.29 | 63.71 | 68.04 | 68.10 | 66.06 | 68.01 | 65.71 | 72.97 | 66.36 | 70.04 | 70.51 | 70.53 | 65.43 |
-| Source | - | 48.68 | 49.79 | 48.97 | 67.69 | 61.82 | 70.93 | 66.18 | 61.39 | 61.37 | 45.29 | 75.99 | 51.89 | 65.78 | 68.71 | 66.14 | 60.71 |
-| TENT* | TTA | 41.93 | 47.19 | 49.22 | 64.80 | 64.33 | 68.05 | 63.86 | 61.42 | 62.12 | 51.65 | 73.92 | 50.80 | 67.47 | 69.03 | 66.80 | 60.17 |
-| SAR | TTA | 48.08 | 49.07 | 48.83 | 66.25 | 63.02 | 70.39 | 66.27 | 59.53 | 60.31 | 46.28 | 75.62 | 51.23 | 65.47 | 68.47 | 66.14 | 60.33 |
-| CoTTA | CTTA | 56.89 | 57.26 | 56.92 | 66.39 | 65.16 | 68.11 | 66.52 | 60.23 | 62.46 | 50.36 | 73.47 | 50.02 | 66.04 | 67.85 | 67.06 | 62.31 |
-| EATA | CTTA | 48.67 | 49.80 | 49.16 | 67.68 | 62.02 | 70.92 | 66.13 | 61.57 | 61.46 | 45.17 | 76.02 | 51.95 | 65.90 | 68.71 | 66.22 | 60.76 |
-| MMTTA* | MM-TTA | 49.82 | 49.90 | 49.85 | 49.94 | 49.86 | 49.92 | 49.89 | 49.97 | 49.94 | 49.79 | 50.54 | 50.01 | 54.33 | 56.20 | 59.99 | 51.33 |
-| READ | MM-TTA | 52.00 | 52.28 | 51.02 | 62.43 | 64.47 | 66.04 | 65.38 | 61.39 | 65.10 | 62.96 | 75.14 | 53.59 | 68.89 | 69.95 | 68.20 | 62.59 |
-| MDAA | MM-CTTA | 70.55 | 70.43 | 70.08 | 71.81 | 70.30 | 72.19 | 70.88 | 69.38 | 69.35 | 67.48 | 73.16 | 62.07 | 68.63 | 67.98 | 65.21 | 69.30 |
-| Method | Type | Gauss. | Shot. | Impul. | Defoc. | Glass. | Motion. | Zoom. | Snow | Frost | Fog | Bright. | Cont. | Elastic. | Pixel. | Jpeg | Avg. |
-| Source | - | 53.05 | 52.91 | 52.98 | 57.19 | 57.37 | 58.37 | 57.45 | 56.37 | 56.47 | 55.41 | 59.19 | 53.75 | 57.19 | 56.44 | 57.33 | 56.10 |
-| TENT* | TTA | 53.27 | 52.76 | 52.00 | 54.58 | 54.35 | 55.04 | 54.86 | 52.59 | 52.81 | 53.11 | 53.50 | 50.80 | 53.03 | 52.31 | 52.14 | 53.14 |
-| SAR | TTA | 53.14 | 53.29 | 53.21 | 56.95 | 56.91 | 58.41 | 57.47 | 56.10 | 56.78 | 56.34 | 58.35 | 53.98 | 57.27 | 55.93 | 56.48 | 56.04 |
-| CoTTA | CTTA | 53.03 | 53.08 | 53.03 | 52.92 | 52.28 | 51.84 | 51.41 | 50.83 | 50.08 | 49.66 | 49.22 | 48.85 | 48.30 | 47.62 | 47.20 | 50.62 |
-| EATA | CTTA | 53.74 | 53.67 | 53.68 | 57.20 | 57.26 | 58.53 | 57.93 | 56.38 | 56.67 | 56.23 | 59.04 | 53.63 | 58.19 | 57.36 | 57.48 | 56.47 |
-| READ | MM-TTA | 53.77 | 54.26 | 54.28 | 58.04 | 58.00 | 59.09 | 58.84 | 57.43 | 58.18 | 58.12 | 59.38 | 55.99 | 58.30 | 57.51 | 57.91 | 57.27 |
-| MDAA | MM-CTTA | 55.63 | 55.91 | 55.88 | 57.50 | 57.77 | 58.27 | 57.84 | 57.05 | 56.72 | 56.77 | 58.67 | 55.78 | 57.55 | 56.87 | 57.84 | 57.07 |
-| Source | - | 53.01 | 52.88 | 52.97 | 57.20 | 57.42 | 58.41 | 57.49 | 56.37 | 56.49 | 55.43 | 59.16 | 53.74 | 57.19 | 56.42 | 57.29 | 56.10 |
-| TENT* | TTA | 50.48 | 50.50 | 50.50 | 53.09 | 53.19 | 53.70 | 53.18 | 52.31 | 53.47 | 53.57 | 55.40 | 52.22 | 56.23 | 56.74 | 56.99 | 53.44 |
-| SAR | TTA | 53.02 | 52.95 | 52.71 | 56.78 | 56.63 | 58.18 | 57.38 | 55.97 | 56.69 | 56.45 | 58.46 | 54.03 | 57.41 | 56.53 | 57.16 | 56.02 |
-| CoTTA | CTTA | 48.10 | 48.51 | 49.13 | 49.90 | 50.47 | 50.93 | 51.67 | 52.14 | 53.17 | 53.19 | 55.86 | 53.21 | 58.02 | 56.60 | 56.71 | 52.51 |
-| EATA | CTTA | 53.69 | 53.66 | 53.61 | 57.26 | 57.34 | 58.46 | 57.84 | 56.35 | 56.82 | 56.72 | 59.21 | 54.10 | 58.28 | 57.39 | 57.54 | 56.55 |
-| READ | MM-TTA | 55.33 | 55.34 | 54.93 | 58.16 | 57.88 | 58.82 | 58.50 | 57.24 | 57.94 | 57.86 | 59.65 | 55.67 | 58.79 | 57.87 | 58.10 | 57.47 |
-| MDAA | MM-CTTA | 55.85 | 55.87 | 55.94 | 57.40 | 57.85 | 58.25 | 57.95 | 57.02 | 56.89 | 56.68 | 58.66 | 55.84 | 57.58 | 56.92 | 58.00 | 57.11 |
-D.3 Dynamic threshold update
-The threshold we used in DSM is a fixed number. In this section we attempt to update in a dynamic way during the adaptation. We define the threshold for classifier in time as
+C.3. Dynamic threshold update
+The threshold we used in DLFM is a fixed number. In this section we attempt to update in a dynamic way during the adaptation. We define the threshold for classifier in time as
 where is the learning rate, is the initial threshold and is the batch size.
 is calculated to reflect the gap between Leader and follower .
-The original intention of this design is to adjust the size of the threshold according to the change of , so as to eliminate the statistical bias of different domains.
-However, as shown in Table 14, such attempt is not only achieve lower performance while needs more variables to be memorized.
-So only the fixed threshold is used in the formal method.
+The original intention is to adjust the threshold size according to changes to eliminate statistical bias across domains.
+However, as shown in Table 15, this approach achieves lower performance while requiring more variables to be memorized.
+Therefore, only the fixed threshold is used in the formal method.
 | KS-video | VGG-audio | |||
@@
 | 0.2 | 62.93 | 69.33 | 34.41 | 34.31 |
-| KS-video | ||
-| 0.9 | 64.29 | 69.29 |
-| 0.8 | 64.88 | 69.29 |
-| 0.7 | 65.43 | 69.30 |
-| 0.6 | 65.20 | 68.81 |
-| 0.5 | 35.37 | 37.37 |
-D.4 Label weight distribution
-The weight assignment in SPS follows the formula .
-In this section we make a toy experiment to explore whether such assignment is reasonable.
-We take KS-video as example which use top 2 classes for reconstruction.
-Table 14 show the results on different weight assignment.
-It can be seen that the performance peaks at 0.7, corresponding to the assignment in main text.
-Therefore the assignment of weights in the main text is justified.
+Appendix D Proofs of Theorems
+Proof of Theorem 1. known the optimal problem in Eqn.3.3 can be further written as:
+Note above equation as , derive for as
+Therefore